### v2

# Matbench Discovery A framework to evaluate machine learning crystal stability predictions

Abstract
Matbench Discovery simulates using machine learning (ML) energy models as pre-filters to DFT in a high-throughput search for stable inorganic crystals. We address the disconnect between (i) thermodynamic stability and formation energy and (ii) in-domain vs out-of-distribution performance. Alongside this paper, we publish a Python package to aid with future model submissions and a growing online leaderboard with further insights into trade-offs between various performance metrics. To answer the question of which ML methodology performs best at materials discovery, our initial release includes random forests, graph neural networks (GNN), one-shot predictors, iterative Bayesian optimizers and universal interatomic potentials (UIP). Ranked best-to-worst by their test set F1 score on thermodynamic stability prediction, we find MACE > CHGNet > M3GNet > ALIGNN > MEGNet > CGCNN > CGCNN+P > Wrenformer > BOWSR > Voronoi fingerprints with random forest. The top 3 models are UIPs which we establish to be by far the winning methodology for ML-guided materials discovery, achieving F1 scores of 0.6 for crystal stability classification and discovery acceleration factors (DAF) of up to 5x on the first 10k most stable predictions compared to dummy selection. We also highlight a disconnect between commonly used regression metrics and more task-relevant classification metrics. Accurate regressors are susceptible to unexpectedly high false-positive rates if those accurate predictions lie close to the decision boundary at above the convex hull where most materials are. Finally, we share valuable insights for maintainers of high throughput materials databases by demonstrating that universal potentials have matured enough to triage and determine compute budget allocation in future expansion efforts.
1 Department of Physics, University of Cambridge, UK
2 Lawrence Berkeley National Laboratory, USA
3 Federal Institute of Materials Research and Testing (BAM), Germany
4 Molecular Foundry, Lawrence Berkeley National Laboratory, USA
5 Department of Materials Science and Engineering, University of California - Berkeley, USA
1 Introduction
Material science can be viewed as a combinatorial problem of mixing and arranging different atoms to leverage the complex range of properties that emerge. To date, combinations have been tested experimentally [1, 2] and have been simulated [3, 4, 5, 6]. [7] identified possible quaternary materials allowed by electronegativity and charge-balancing rules. The space of quinternaries and higher is even less explored, leaving vast numbers of potentially useful materials to be discovered. The discovery of new materials is a key driver of technological progress and lies on the path to more efficient solar cells, lighter and longer-lived batteries, smaller and more efficient transistor gates just to name a few. In light of our sustainability goals, these advances cannot come fast enough. Any speed-up new discovery methods might yield should be leveraged to their fullest extent.
Despite significant advances in empirical, theoretical and computational materials science, discovering new materials still requires complex calculations, labor-intensive trial-and-error experimentation, and often happens fortuitously rather than through rational design. Machine learning (ML) methods efficiently extract and distill trends from huge datasets, and can handle high dimensionality, multiple objectives [8], uncertainty [9, 10, 11], and noisy or sparse data [12, 13], making them powerful additions to the computational materials science tool set.
ML models are less accurate and reliable but orders of magnitude faster than ab initio simulation. This makes them most suitable for use in high-throughput (HT) searches to triage more expensive, higher-fidelity simulation methods. The use of neural networks for learning the Kohn-Sham density-functional theory (DFT) potential energy surface (PES) can be traced as far back as [14]. This work kicked off rapid advances and significant efforts to fit ever more sophisticated ML models to known samples of the PES. Initially, most of these models were trained and deployed as interatomic potentials (also known as force fields) to study known materials of interest, a workflow that requires curating custom training data for each new system of interest [15, 16]. As larger and more diverse datasets have emerged from initiatives like the Materials Project (MP) [3], AFLOW [5] or the Open Quantum Materials Database (OQMD) [4], researchers have begun to train so-called universal models that cover 90 or more of the most-application relevant elements in the periodic table. This opens up the prospect of ML-guided materials discovery to increase the hit rate of stable crystals and speed up DFT- and expert-driven searches.
Progress in ML for materials is often measured according to performance on standard benchmark data sets. As ML models have grown in complexity and applicability, benchmark datasets need to grow with them to accurately measure their usefulness. However, due the the rapid pace of the field and the variety of possible approaches for framing the discovery problem, no large-scale benchmark yet exists for measuring the ability of ML to accelerate materials discovery. As a result, it is unclear which methodologies or models are best suited for this task. Recent works focusing on prospective computational materials discovery have proposed strategies based on one-shot coordinate free predictors [10], iterative Bayesian optimizers [17], and universal interatomic potentials (UIP) [18, 19, 20]. These papers deploy their respective model on specific systems and custom datasets to highlight certain strengths but have yet to be compared in a systematic standardized manner that allows them to be ranked by their ability to accelerate materials discovery. Our work aims to identify the state-of-the-art (SOTA) model by proposing a novel evaluation framework that closely simulates a real-world discovery campaign guided by ML models.
Specifically, we designed a benchmark task that tackles three central challenges. We believe these challenges to be requirements when seeking to justify experimental validation of ML predictions:
1.
Mimic Real Applications: Idealized and overly simplified benchmarks can fail to reflect the real-world challenges a model faces when used in an actual discovery campaign. This can result in a disconnect between benchmark metrics and real-world performance. Possible reasons for this include choosing the wrong target [13] or picking an unrepresentative train/test split [21, 22]. In the case of materials discovery, formation energies - although widely used as regression targets - do not indicate thermodynamic stability. Distance to the convex hull spanned by competing phases in the same chemical system is a much more informative indicator of thermodynamic stability. Moreover, ML models relying on relaxed crystal structures as input render any discovery pipeline circular since obtaining relaxed structures requires computationally expensive DFT simulations, thereby depending on the very process we intend to accelerate.
2.
Opportunity cost: Accurate regressors are susceptible to unexpectedly high false-positive rates if those accurate predictions lie close to the decision boundary. Looking purely at global metrics like MAE, RMSE and can give practitioners a false sense of security about their model’s reliability. Failed experiments incur a high opportunity cost by wasting lab time and resources.
3.
Scalability: Future discovery efforts are likely to encounter large data regimes. Small benchmarks can lack chemical diversity, obfuscate poor scaling relations or poor out-of-distribution (OOD) performance (see fig. 10 for a UMAP projection of train and test sets). For instance, random forests achieve excellent performance on small datasets but are typically outperformed by neural networks on large datasets due to the benefits of representation learning [23].
Two prior efforts worth highlighting that have partially addressed the above challenges are Matbench [24] and the Open Catalyst Project (OCP) [25].
By providing a standardized collection of 13 datasets ranging in size from 300 to 132,000 samples from both DFT and experimental sources, Matbench addresses the scalability challenge, highlighting how model performance changes as a function of data regime. Matbench helped focus the field of ML for materials, increase comparability across papers and provide a quantitative measure of progress in the field. Importantly, all tasks were exclusively concerned with the properties of known materials. We believe a task that simulates a materials discovery campaign by requiring materials stability prediction from unrelaxed structures to be a missing piece here.
OCP is a large-scale initiative to discover substrate-adsorbate combinations that catalyze key industrial reactions that process said adsorbates into more useful products. The OCP has released two data sets thus far, OCP20 [25] and OCP22 [26], for training and benchmarking ML models. OCP certainly addressed challenge 1 of closely mimicking a real-world problem. They have recently shown that despite not reaching their target accuracy to entirely replace DFT, using ML in conjunction with confirmatory DFT calculations dramatically speeds up their combinatorial screening workflow [27].
We believe that addressing these three challenges will enable future ML-guided discovery efforts to expand materials databases to confidently select appropriate models and methodology. Our findings show that UIPs outperform all other methodologies we tested both in accuracy and robustness.
1.1 Evaluation Framework for Materials Discovery
We propose a novel evaluation framework that places no constraints on the type of data a model trains on as long as it would be available to a practitioner conducting a real materials discovery campaign. We only enforce that at test time, all models must make predictions on the convex hull distance of the relaxed structure with only the unrelaxed structure as input. This setup avoids circularity in the discovery process, as unrelaxed structures can be cheaply enumerated through elemental substitution methodologies and do not contain information inaccessible in a prospective discovery campaign. We choose to focus on the relaxed structure’s convex hull distance as a measure of thermodynamic stability rather than the formation energy as it informs the decision on whether to pursue a potential candidate crystal. This decision was also motivated by [13] who found even composition-only models capable of predicting DFT formation energies with useful accuracy. However, when tasking those same models with predicting decomposition enthalpy, performance deteriorated sharply. This insight meant that ML models are much less useful than DFT for discovering new inorganic solids. Moreover, the qualitative leap in performance from Roost [23], the best compositional model benchmarked, to CGCNN [28], the single structural model they tested, highlights that structure plays a crucial role in determining material stability. Hence our decision to restrict test-time model input to unrelaxed structures to avoid the above-mentioned circularity and measure true prospective utility while still allowing for subtle atomic-configuration-informed energy estimates. We recognize that our ground truth is ignorant of entropic stabilization and metastable states. While these factors influence experimental synthesizability, we have to neglect them when simulating an HT search since the convex hull distance as predicted by zero-temperature DFT is the best proxy of true crystal stability available for large datasets.
Standard practice in ML benchmarks is to hold all variables fixed – most importantly the training data – and vary only the model architecture to isolate architectural effects on the performance. We deliberately deviate from this practice due to diverging objectives from common ML benchmarks. Our goal is to identify the best methodology for accelerating materials discovery. What kind of training data a model can ingest is part of its methodology. Unlike energy-only models, UIPs benefit from the additional training data provided by the forces and stresses recorded in DFT relaxations. This allows them to learn a fundamentally higher fidelity model of the physical interactions between ions. That is a genuine advantage of the architecture and something any benchmark aiming to identify the optimal methodology for materials discovery must reflect, as ours does. As a result, our benchmark contains models trained on varying data sets.
We define the Materials Project (MP) [3] database release (v2022.10.28 as the maximum allowed training set for any valid model submission. Any subset of the energies, forces, stresses, magnetic moments or any other recorded DFT data are valid training targets. All of these would be available to a practitioner performing a real materials discovery campaign and hence are permitted for any model submission. Models may train on the complete set of relaxation frames, or any subset thereof such as the final relaxed structures only. Our test set consists of the unrelaxed structures in the WBM dataset [29]. Their target values are the PBE formation energies of the corresponding DFT-relaxed structures.
1.1.1 Materials Project Training Set
The Materials Project is a well-known effort to calculate the properties of all inorganic materials using high-throughput ab initio methods. Seeded from a subset of the Inorganic Crystal Structure Database (ICSD) [30], the initial release of the database consisted of 9 k crystals. At the time of writing, the Materials Project database [3] has grown to 154 k crystals, covering diverse chemistries and providing relaxed and initial structures as well as the relaxation trajectory for every entry.
Our benchmark defines the training set as all data available from the v2022.10.28 MP release. We recorded a snapshot of energies, forces, stresses and magnetic moments for all MP ionic steps on 2023-03-15 as the canonical training set for Matbench Discovery, and provide convenience functions through our Python package for easily feeding that data into future model submissions to our benchmark.
The flexibility in specifying the maximum allowed data set is intended to allow authors to experiment with and exploit the large variety of available data. This choice is motivated by two factors. First, whilst it may appear that models that train on multiple snapshots containing energies, forces and stresses are getting more training data than models that are only trained or able to be trained on the energies of the relaxed structures the critical factor is that all this additional data was generated as a byproduct of the workflow to produce the relaxed structures i.e. all models have been trained from the same total cost of training data acquisition. If some architectures or approaches can leverage more of this byproduct data to make improved predictions this is a fair comparison between the two models. This approach diverges philosophically from other benchmarks such as the OCP and Matbench where it has been more common to subcategorize different models and look at multiple alternative tasks (e.g. composition-only vs structure-available in Matbench or IS2RS, IS2RE, S2EF in OCP) and do not make direct comparisons of this manner. Second, recent work in the space from [31, 32] has claimed that much of the data in large databases like MP are redundant and that models can be trained more effectively by taking a subset of these large data pools. We believe that from a systems-level perspective identification of novel cleaning and or active-learning strategies that do not leak signal from the test set are equally as important as architectural improvements as they can lead to similar improvements in performance, particularly given the prevalence of errors in high-throughput DFT that can disrupt learning of the PES. Consequently, such strategies where they lead to improved performance should be able to be recognized within the benchmark. We encourage submissions to submit ablation studies showing how different system-level choices affect performance.
We highlight for example data sets that are valid within the rules of the benchmark that take advantage of these freedoms. The first is the MPF.2021.2.8 data set [18] curated to train the M3GNet model which takes a subset of materials from the v2021.02.08 MP release. Rather than taking every ionic step from the relaxation trajectory this data set opts to select only the initial, final and one intermediate structure for each material to avoid biasing the data set towards examples where more ionic steps were needed to relax the structure. Additionally, the curators of the MPF.2021.2.8 data set down-sampled the v2021.02.08 release significantly to select a subset of calculations that they believed to be most self-consistent. The MPF.2021.2.8 is a proper subset of the training data as no materials were deprecated between the v2021.02.08 and v2022.10.28 database releases. The second data set we highlight, with which several of the UIP models have been trained, is the MPtrj dataset [19]. This data set was curated from the earlier v2021.11.10 MP release. The MPtrj dataset is a proper subset of the allowed training data but several potentially anomalous examples from within MP were cleaned out of the data set before the frames were subsampled to remove. It is worth noting that the v2022.10.28 release contains a small number of additional Perovskite structures not found in MPtrj that could be added to the training set within the scope of the benchmark.
We note that the v2023.11.1 deprecated a large number of calculations so data queried from subsequent database releases is not considered valid for this benchmark.
1.1.2 WBM Test Set
The WBM data set [29] consists of 257 k structures generated via chemical similarity-based elemental substitution of MP source structures followed by DFT relaxation and calculating each crystal’s convex hull distance. Which element replaces an existing one in a given source structure was determined by random sampling according to the weights in a chemical similarity matrix data-mined from the ICSD [33].
The WBM authors performed 5 iterations of this substitution process (we refer to these as batches). After each step, the newly generated structures found to be thermodynamically stable after DFT relaxation flow back into the source pool to partake in the next round of substitution. This split of the data into batches of increasing substitution count is a unique and compelling feature of the test set as it allows out-of-distribution (OOD) testing by seeing if model performance degrades with substitution count. A higher number of elemental substitutions on average carries the structure further away from the region of material space covered by the MP training set (see figs. 8 and 10 for details). Whilst this batch information makes the WBM data set an exceptionally useful data source for examining the extrapolation performance of ML models, our evaluation primarily looks at metrics that consider all batches as a single test set.
Throughout this work, we define stability as being on or below the convex hull of the MP training set (). 42k out of 257k materials in WBM satisfy this criterion. Of these, 33k are unique prototypes, meaning they have no matching structure prototype in MP nor another higher-energy duplicate prototype in WBM. Our code treats the stability threshold as a dynamic parameter, allowing for future model comparisons at different thresholds. This may reveal systematic differences in stability calibration, as hinted by the different TPR/TNR tradeoffs observed for CHGNet and M3GNet in table 3. A looser stability threshold (e.g. ) might benefit a model like M3GNet which is more conservative in its stability predictions, whereas a stricter threshold may benefit a model like CHGNet that is more optimistic in its stability predictions. For initial analysis in this direction, see fig. 3 in the SI.
As WBM explores regions of materials space not well sampled by MP, many of the discovered materials that lie below MP’s convex hull are not stable relative to each other. Of the 42k that lie below the MP convex hull less than half, or around 20k, remain on the joint MP+WBM convex hull. This observation suggests that many WBM structures are repeated samples in the same new chemical spaces. It also highlights a critical aspect of this benchmark in that we purposely operate on an incomplete convex hull. Only current knowledge of competing points on the PES is accessible to a real discovery campaign and our metrics are designed to reflect this.
1.2 Models
To test a wide variety of methodologies proposed for learning the potential energy landscape, our initial benchmark release includes 10 models.
1.
MACE [20] (UIP-GNN) - MACE builds upon the recent advances [34, 35] in equivariant neural network architectures by proposing an approach to computing high-body-order features efficiently via Atomic Cluster Expansion [36]. Unlike the other UIP models considered MACE was primarily developed for molecular dynamics of single material systems and not the universal use case studied here. It is the only -equivariant model we tested.
2.
CHGNet [19] (UIP-GNN) - CHGNet is a UIP for charge-informed atomistic modeling. Its distinguishing feature is that it was trained to predict magnetic moments on top of energies, forces and stresses in the MPTrj dataset consisting of relaxation trajectories for 1.5 million MP structures (see section A.10 for detailed analysis of this dataset). By modeling magnetic moments, CHGNet learns to accurately represent the orbital occupancy of electrons which allows it to predict both atomic and electronic degrees of freedom.
3.
M3GNet [18] (UIP-GNN) - M3GNet is a GNN-based universal interatomic potential (UIP) (as in covering the full periodic table) for materials trained on up to 3-body interactions in the initial, middle and final frame of MP DFT relaxations. The model takes the unrelaxed input and emulates structure relaxation before predicting energy for the pseudo-relaxed structure.
4.
ALIGNN [37] (GNN) - The Atomistic Line Graph Neural Network (ALIGNN) is a message passing GNN architecture that takes as input both the interatomic bond graph and a line graph corresponding to 3-body bond angles. The ALIGNN architecture involves a global pooling operation which means that it is ill-suited to force-field applications. To address this the ALIGNN-FF model was later introduced without global pooling [38].
5.
MEGNet [39] (GNN) - MatErials Graph Network is another GNN similar to CGCNN for material properties of relaxed structures that also updates the edge and global features (like pressure, temperature, entropy) in its message passing operation. This work showed that learned element embeddings encode periodic chemical trends and can be transfer-learned from large data sets (formation energies) to predictions on small data properties (band gaps, elastic moduli).
6.
CGCNN [28] (GNN) - The Crystal Graph Convolutional Neural Network (CGCNN) was the first neural network model to directly learn 8 different DFT-computed material properties from a graph representing the atoms and bonds in a periodic crystal. CGCNN was among the first to show that just like in other areas of ML, given large enough training sets, neural networks can learn embeddings that outperform human-engineered structure features directly from the data.
7.
CGCNN+P [40] (GNN) - This work proposes simple, physically motivated structure perturbations to augment stock CGCNN’s training data of relaxed structures with structures resembling unrelaxed ones but mapped to the same DFT final energy. Here we chose , meaning the training set is augmented with 5 random perturbations of each relaxed MP structure mapped to the same target energy.
In contrast to all other structure-based GNNs considered in this benchmark, CGCNN+P is not attempting to learn the Born-Oppenheimer potential energy surface. The model is instead taught the PES as a step-function that maps each valley to its local minimum. The idea is that during testing on unrelaxed structures, the model will predict the energy of the nearest basin in the PES. The authors confirm this by demonstrating a lowering of the energy error on unrelaxed structures. -
8.
Wrenformer (Transformer) - For this benchmark, we introduce Wrenformer which is a variation on the coordinate-free Wren model [10] constructed using standard QKV-self-attention blocks [41] in place of message-passing layers. This architectural adaptation reduces the memory usage allowing the architecture to scale to structures with greater than 16 Wyckoff positions.
Like its predecessor, Wrenformer is a fast coordinate-free model aimed at accelerating screening campaigns where even the unrelaxed structure is a priori unknown. The key idea is that by training on the coordinate anonymized Wyckoff positions (symmetry-related positions in the crystal structure), the model learns to distinguish polymorphs while maintaining discrete and computationally enumerable inputs. The central methodological benefit of an enumerable input is that it allows users to predict the energy of all possible combinations of spacegroup and Wyckoff positions for a given composition and maximum unit cell size. The lowest-ranked prototypes can then be fed into downstream analysis or modeling. -
9.
BOWSR [17] (BO-GNN) - BOWSR combines a symmetry-constrained Bayesian optimizer (BO) with a surrogate energy model to perform an iterative exploration-exploitation-based search of the potential energy landscape. Here we use MEGNet [39] for the energy model as proposed in the original work. The high sample count needed to explore the PES with BO makes this by far the most expensive model tested.
10.
Voronoi RF [42] (Fingerprint) - A random forest trained to map a combination of composition-based Magpie features [43] and structure-based relaxation-robust Voronoi tessellation features (effective coordination numbers, structural heterogeneity, local environment properties, …) to DFT formation energies.
This fingerprint-based model predates most deep learning for materials but significantly improved over earlier fingerprint-based methods such as the Coulomb matrix [44] and partial radial distribution function features [45]. It serves as a baseline model to see how much value the learned featurization of deep learning models can extract from the increasingly large corpus of available training data.
2 Results
Table 1 shows performance metrics for all models included in the initial release of Matbench Discovery. MACE takes the top spot and emerges as the current SOTA for ML-guided materials discovery. It outperforms the 9 other models on all 6 classification metrics but is slightly overtaken by CHGNet in 2 of 3 regression metrics, and RMSE, which are sensitive to outliers. The discovery acceleration factor (DAF) measures how many more stable structures a model found compared to the dummy discovery rate achieved by randomly selecting test set crystals, formally the DAF is the ratio of the precision to the prevalence. The maximum possible DAF is the inverse of the prevalence or dummy discovery rate which on our dataset is . Thus the current SOTA of 3.85 achieved by MACE leaves room for improvement. However, evaluating each model on the much smaller subset of its 10k most stable predictions as shown in table 3, CHGNet takes the lead in all 9 metrics, in particular achieving an impressive DAF of 5.29 which approaches optimal performance.
We find a large performance gap between models that make one-shot predictions directly from unrelaxed inputs (MEGNet, Wrenformer, CGCNN, CGCNN+P, ALIGNN, Voronoi RF) compared to UIPs that learn from force and stress labels to emulate DFT relaxation (CHGNet, M3GNet, MACE). While the F1 scores and DAFs of force-free models are surprisingly less affected, their regression metrics such as and RMSE are significantly worse. Of the force-free models, only ALIGNN, BOWSR and CGCNN+P achieve positive predictive power . Negative means model predictions explain the observed variation in the data less than simply predicting the test set mean. In other words, these models are not predictive in a global sense (across the full dataset range). However, even models with negative can be locally good in the positive and negative tails of the test set distribution. They suffer most in the mode of the distribution near the stability threshold of above the hull. This reveals an important shortcoming of as a metric for classification tasks like stability prediction.
The results for M3GNet depart from the trend that F1 is broadly rank-correlated with the other classification metrics. M3GNet achieves a high true positive rate () but a notably lower true negative rate () than CHGNet and ALIGNN with the closest F1 scores. Figure 2 provides a visual explanation for this observation. M3GNet has a lower rolling mean of the absolute errors (rolling MAE) as a function of hull distance for materials above the convex hull (see right half of plot) but incurs comparably large errors for materials below the hull (left half of plot). Since , lower error for energies above the hull increases both TN and decreases FP, resulting in the high TPR values observed.
The reason CGCNN+P achieves better regression metrics than CGCNN but is still worse as a classifier becomes apparent from fig. 7 by noting that the CGCNN+P histogram is more sharply peaked at the 0 hull distance stability threshold. This causes even small errors in the predicted convex hull distance to be large enough to invert a classification. Again, this is evidence to choose carefully which metrics to optimize. Regression metrics are far more prevalent when evaluating energy predictions. However, our benchmark treats energy predictions as merely means to an end to classify compound stability. Improvements in regression accuracy are of limited use to materials discovery in their own right unless they also improve classification accuracy. Our results demonstrate that this is not a given.
Figure 1 has models rank materials by model-predicted hull distance from most to least stable; materials furthest below the known hull at the top, materials right on the hull at the bottom. For each model, we iterate through that list and calculate at each step the precision and recall of correctly identified stable materials. This simulates exactly how these models would be used in a prospective materials discovery campaign and reveals how a model’s performance changes as a function of the discovery campaign length. As a practitioner, you have a certain amount of resources available to validate model predictions. These curves allow you to read off the best model given these conditions and based on the optimal trade-off between fewer false positives (precision) or fewer negatives (recall) for the discovery task at hand. In this case, it so happens that CHGNet achieves the highest precision and recall at any number of screened materials.
A line terminates when a model believes there are no more materials in the WBM test set below the MP convex hull. The dashed vertical line shows the actual number of stable structures in our test set. All models are biased towards stability to some degree as they all overestimate this number, most of all BOWSR by 133%, least of all MEGNet by 30%. This is only a problem in practice for exhaustive discovery campaigns that validate all stable predictions from a model. More frequently, model predictions will be ranked most-to-least stable and validation stops after some pre-determined compute budget is spent, say, 10k DFT relaxations. In that case, the concentration of false positive predictions that naturally accumulates near the less stable end of the candidate list can be avoided with no harm to the campaign’s overall discovery rate (see table 3 a similar metrics table considering only the 10k materials predicted by each model to be furthest below the known convex hull).
The diagonal Optimal Recall line would be achieved if a model never made a false negative prediction and stopped predicting stable crystals exactly when the true number of stable materials is reached. Zooming in on the top-left corner of the precision plot, we observe that CHGNet is the only model without a sudden drop in precision right at the start of the discovery campaign. This means CHGNet is the only model whose first few hundred most stable materials are largely actually stable. MACE in particular suffers from several initial failure cases. MACE underpredicts the energy of these unstable materials by several eV/atom (see fig. 4), resulting in MACE’s initial precision starting at 0 before abruptly climbing to 0.8. We experienced this behavior with multiple independently trained MACE variants; even more so with a checkpoint we received directly from the MACE authors which was trained on the M3GNet training set. The results shown here are from a superior MACE we trained ourselves on the much larger MPtrj dataset. Even M3GNet, the other universal potential and 2nd best model, exhibits the early-on precision drop. As a result, CHGNet has a strong lead over M3GNet until reaching 3k screened materials. From there, CHGNet and M3GNet slowly converge until they almost tie at a precision of 0.52 after 56k screened materials. At that point, CHGNet’s list of stable predictions is exhausted while M3GNet continues, dropping in precision to 0.45 at 76 k, attributable to many false positives near the end of the list of stable predictions.
All force-free models exhibit a much worse case of early-on precision drop, falling to 0.6 or less in the first 5k screened materials. Many of these models (all except BOWSR, Wrenformer and Voronoi RF) display an interesting hook shape in their cumulative precision, recovering again slightly in the middle of the simulated campaign between 5k and up to 30k before dropping again until the end.
Figure 2 provides a visual representation of the reliability of different models based on the rolling mean absolute error (MAE) of model-predicted hull distances as a function of DFT distance to the Materials Project (MP) convex hull. The red-shaded area, which we coin the ‘triangle of peril’, emphasizes the zone where the average model error surpasses the distance to the stability threshold at 0 eV. As long as the rolling MAE remains within this triangle, the model is most susceptible to misclassifying structures. Because the average error is larger than the distance to the classification threshold at 0, it is large enough to flip a correct classification into an incorrect one (if the error happens to point toward the stability threshold).
The sooner a model leaves the triangle on the left side, the less likely it is to incorrectly predict stable structures as unstable, thereby reducing false negatives. On the right side, early exits result in a lower likelihood of predicting unstable structures as stable, leading to fewer false positives.
On the left side, CHGNet exits the triangle first followed closely by ALIGNN. We can expect fewer false negatives from CHGNet than from other models. On the right side, CHGNet and M3GNet exit simultaneously, at a remarkably low error, followed shortly after by MACE at . All 3 UIPs therefore produce fewer false positives than other models. For highly unstable structures, first CHGNet and later M3GNet are both overtaken by Wrenformer for highly unstable structures towards the right end of the plot.
Overall, this visualization underscores the importance of considering a model’s average error relative to the distance to the stability threshold when evaluating its performance as a material stability predictor. Low overall hull distance error can be a misleading metric if that error is not particularly low at the decision boundary.
The imbalance between the left and right half of the plot shows that models are more prone to false negative predictions, even for very stable materials far below the known hull, than predicting very unstable materials as stable.
As the convex hull becomes more thoroughly sampled by future discovery, the fraction of unknown stable structures decreases, naturally leading to less enriched future test sets. This has several implications, including allowing for higher maximum DAFs but also skewing the hull distance distribution towards positive values (i.e. the right half of fig. 2) as there are fewer undersampled chemical spaces. Consequently, to have accurate stability classifications, model accuracy will need to improve concomitantly. For fig. 2, this means models need to be much more accurate to exit the shaded triangle in the left half of the plot.
3 Discussion
We have demonstrated the effectiveness of ML-based triage in HT materials discovery and posit that the benefits of including ML in discovery workflows now clearly outweigh the costs. LABEL:{tab:metrics-table-uniq-protos} shows in a realistic benchmark scenario that several models achieve a discovery acceleration greater than 2.5 across the whole dataset and up to 5 when considering only the 10k most stable predictions from each model (table 3). When starting this project, we were unsure which is the most promising ML methodology for HT discovery. Our findings demonstrate a clear superiority in the accuracy and extrapolation performance of UIPs like CHGNet, M3GNet and MACE. Modeling forces enables these models to chart a path through atomic configuration space closer to the DFT-relaxed structure from where a more informed final energy prediction is possible. The consistently linear log-log learning curves observed in related literature [46] suggest that further decreases in the error of UIPs can be readily unlocked with increased training data.
Despite its fast pace of progress, the field of universal ML potentials is still in its early stages. To realize the full potential of these models, we will need to pivot significant resources to generate large quantities of higher-than-PBE fidelity training data. That said, we believe our work (e.g. fig. 8) and others [27] have caught early glimpses that sufficiently large training sets allow us to narrow the performance gap between in-domain and out-of-domain performance. This should further increase the discovery acceleration factor, resulting in a virtuous cycle between ML models accelerating the creation of new training data while becoming more accurate and even stronger accelerators after retraining. Meanwhile, the setup cost and learning curve of ML potentials are steadily decreasing, likely with a much lower floor than DFT given the intricacies of HT ab-initio simulation. Combined with their already strong performance, this suggests that UIPs combined with data accumulation over the next few years may unlock our way to a comprehensive yet cheap map of the PES that truly commoditizes HT materials discovery.
Despite this optimistic outlook, not only has the predictive power UIPs attained for crystal stability over the last 3 years since [13] sobering 2020 analysis yet to enter common knowledge, but many remaining shortcomings invite further research. For example, all models are biased towards stability, overestimating the correct number of stable materials in our test set anywhere between 30% and 133%. Moreover, even the best models in our analysis still exhibit a significant fraction of false positive predictions, even for materials as high as above the hull. This indicates the need to train future models on even more diverse datasets that are less biased towards ground state crystals than the MP training set we used. No such dataset built with MP-compatible DFT settings exists to our knowledge, inviting future efforts in this direction.
Looking beyond mere thermodynamic stability prediction at zero Kelvin for the purpose of materials discovery, future materials science will also require understanding and predicting the properties under varying environmental conditions such as finite temperature and pressure. This is where interatomic potentials can unlock further utility. Their force predictions enable insights into dynamical properties. One open question of particular relevance is the extent to which such models may aid research into the computational prediction of synthesis pathways. Many existing approaches for reaction pathway prediction involve the use of heuristic rules for dealing with the significant added complexity of metastability alongside traditional ground state ab-initio data [47, 48, 49]. These algorithms stand to benefit massively from more efficient estimates of the reaction energy barriers that future UIPs may provide. Provided they reach sufficient accuracy, this may enable timely calculation of approximate reaction barriers [50] and open up a whole new field to high-throughput inquiry.
4 Acknowledgments
J.R. acknowledges support from the German Academic Scholarship Foundation (Studienstiftung). A.A.L. acknowledges support from the Royal Society. A.J. and K.A.P. acknowledge the US Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division under contract no. DE-AC02-05-CH11231 (Materials Project program KC23MP). This work used computational resources provided by the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231.
Our profound gratitude extends to Hai-Chen Wang, Silvana Botti and Miguel A. L. Marques for their valuable contribution in crafting and freely sharing the WBM data set.
We thank Rickard Armiento, Felix A. Faber and Abhijith S. Parackal for helping develop the evaluation procedures for Wren upon which this work builds. We also thank Rokas Elijosius for assisting in the initial implementation of Wrenformer.
We would like to thank Jason Blake Gibson, Shyue Ping Ong, Chi Chen, Tian Xie, Peichen Zhong and Ekin Dogus Cubuk for helpful discussions.
5 Author Contributions
Janosh Riebesell: Methodology, Software, Data Curation, Training and Testing Models, Formal Analysis. Rhys Goodall: Conceptualization, Software, Formal Analysis. Philipp Benner: Software, Training ALIGNN. Yuan Chiang: Training MACE. Bowen Deng: Curating MPtraj dataset and training CHGNet. Alpha Lee: Supervision. Anubhav Jain: Supervision. Kristin Persson: Funding Acquisition.
6 Code availability
We welcome further model submissions to our GitHub repository https://github.com/janosh/matbench-discovery.
7 Data availability
We chose the latest Materials Project (MP) [3] database release (v2022.10.28 at time of writing) as the training set and the WBM dataset [29] available at https://figshare.com/articles/dataset/22715158 as the test set for this benchmark. A snapshot of every ionic step including energies, forces, stresses and magnetic moments in the MP database is available at https://figshare.com/articles/dataset/23713842.
Appendix A Supplementary Information
A.1 Metrics for 10k materials predicted most stable
Unlike table 1 which evaluates model performance on the subset of unique WBM structure prototypes, table 2 includes the full WBM test set of 257k materials. The 44k additional materials in table 2 comprise discarded earlier due to having a matching structure prototype in MP plus materials which are prototype duplicates of another WBM material with lower energy. The most noteworthy difference between the two tables is a drop in DAF for all models in table 2 compared to table 1. MACE for example achieves a DAF of 3.5 on the full test set (table 2) compared to 3.85 on the subset of 215.5k materials with unique prototypes (table 1). This 10% increase is largely due to a 10% decrease in the fraction of materials below the MP convex hull: 15.3% (32,942 out of 215,488) in table 1 vs 16.7% (42,825 out of 256,963) in the full dataset (table 2). Since DAF is the ratio of model’s precision when predicting stability to the prevalence of stable structures, a lower prevalence results in a higher DAF.
While prototypes are non-trivial to match, thus potentially introducing bias into the test set by trying to deduplicate them, we still opted to feature table 1 in the main text since the removal of overlapping prototypes with MP makes it more closely reflect a model’s true ability to extrapolate to out-of-domain materials. Prototypes were matched based on an Aflow-style [51] Wyckoff representation encoding the crystal’s prototype as implemented in aviary.wren.utils.get_aflow_label_from_spglib| [10]. This string is invariant to relaxation (unless the system changes spacegroup during relaxation, usually by gaining symmetry) and will count structures with different lattice parameters as duplicates if both are expected to relax to the same ground state.
A real-world discovery campaign is unlikely to validate all stable predictions from a given model as we did in LABEL:{tab:metrics-table-uniq-protos}. Presumably, it will rank model predictions from most to least stable and follow that list as far as time and compute budget permits. Assuming that increases in compute resources will allow average future discovery campaigns to grow in scope, we believe 10k model validations to be a reasonable scope for average campaigns. This is what table 3 simulates by calculating classification and regression metrics for the 10k test set materials predicted to be most stable by each model. We again show dummy performance in the bottom row. Note that each model is now evaluated on a different slice of the data. However, the bottom row still shows dummy performance across the whole dataset. CHGNet and M3GNet achieve an impressive precision of 0.83 and 0.8, respectively. In concrete terms, this means in a discovery campaign that validates 10 k model predictions from a search pool of 257 k crystals that are chemically dissimilar from the training set and of which 0.167 are stable, CHGNet and M3GNet would deliver 4 stable structures for every 5 predictions validated. We emphasize that in light of the significant resulting increase in stability hit rate, these models are well worth integrating into future materials searches.
A.2 ROC Curves
A material is classified as stable if the predicted is below the stability threshold. Since all models predict (and M3GNet predicted energies are converted to formation energy before stability classification), they are insensitive to changes in the threshold . The receiver operating characteristic (ROC) curve for each model is plotted in fig. 3. M3GNet wins in the area under curve (AUC) with 0.87, coming in 1.34x higher than the worst model, the Voronoi Random Forest. The diagonal ‘No skill’ line shows the performance of a dummy model that randomly ranks material stability.
A.3 Parity Plots
Figure 4 shows that all models do well for materials far below the convex hull (left side of the plot). Performance for materials far above the convex hull is more varied with occasional underpredictions of the energy of materials far above the convex hull (right side). All models suffer most in the mode of the distribution at .
Two models stand out as anomalous to the general trends.
Wrenformer is the only model with a large number of severe energy overpredictions at along the positive axis. We investigate these failure cases in more detail in fig. 5 and find these overpredictions to be dominated by spacegroup 71 with poor representation in the training data.
The other anomalous model is MACE with several severe underpredictions at along the negative axis. We investigated these points for common traits in composition or crystal symmetry but noticed no pattern.
Beyond these MACE outliers visible in the plot, MACE exhibits another rare but reproducible type of failure case, in which the final predicted energy after relaxation is off by several orders of magnitude. The largest ‘derailed’ prediction was eV/atom for wbm-3-31970 (formula HIr). In each case, the MACE relaxation exhausted the maximum number of ionic steps set to 500 and caused a volume implosion from initial cell volumes of hundreds to relaxed cell volumes of tens of cubic Angstrom. Using the checkpoint trained on the M3GNet dataset which we received from the MACE authors, this failure mode occurred for several hundred of the 250k test set crystals. Using the checkpoint we trained ourselves on the MPTrj dataset, it affects only 44 test crystals, suggesting that these holes in the MACE PES can perhaps be fully plugged by further increasing the training set or even changing the loss function. Further analysis is ongoing. Since these derailed values are easily identified in practice when actually performing a prospective discovery campaign, we excluded them from the MACE parity plat and all other downstream analyses.
A.4 Spacegroup prevalence in Wrenformer failure cases
Figure 5 shows 456 + 194 (70%) of the failure cases in the shaded rectangle are two prototypes in spacegroup 71. The occurrence of those same prototypes in the MP training set shows almost no data support for the failing prototypes. This suggests the reason Wrenformer fails so spectacularly on these structures is that it cannot deal with structure prototypes it has not seen at least several hundred examples of in its training data. This suggests that there are stronger limitations on how much the discrete Wyckoff-based representation can extrapolate to new prototypes compared to the smooth local-environment-based inputs to GNN-type models.
A.5 Hull Distance Box plot
BOWSR has the largest median error, while Voronoi RF has the largest IQR. Note that MEGNet and CGCNN are the only models with a positive median. Their hull distance errors are biased towards more frequently predicting thermodynamic instability, explaining why they are closest to correctly predicting the overall number of stable structures in the test set (see fig. 1).
A.6 Classification Histograms using Model-Predicted Energies
Note the CGCNN+P histogram is more strongly peaked than CGCNN’s which agrees better with the actual DFT ground truth distribution of hull distances in our test set. This explains why CGCNN+P performs better as a regressor, but also reveals how it can perform simultaneously worse as a classifier. By moving predictions closer to the stability threshold at above the hull, even small errors are significant enough to tip a classification over the threshold.
A.7 Measuring extrapolation performance from WBM batch robustness
As a reminder, the WBM test set was generated in 5 successive batches, each step applying another element replacement to an MP source structure or a new stable crystal generated in one of the previous replacement rounds. The likelihood by which one element replaces another is governed by ISCD-mined chemical similarity scores for each pair of elements. Naively, one would expect model performance to degrade with increasing batch count, as repeated substitutions should on average ‘diffuse’ deeper into uncharted regions of material space, requiring the model to extrapolate more. We observe this effect for some models much more than others.
Figure 8 shows the rolling MAE as a function of distance to the convex hull for each of the 5 WBM rounds of elemental substitution. These plots show a stronger performance decrease for Wrenformer and Voronoi RF than for UIPs like M3GNet, CHGNet, MACE and even force-less GNNs with larger errors like MEGNet and CGCNN.
MEGNet and CGCNN both incur a higher rolling MAE than Wrenformer across all 5 batches and across most or all of the hull distance range visible in these plots. However, similar to the UIPs, MEGNet and CGCNN exhibit very little degradation for higher batch counts. The fact that higher errors for later batches are specific to Wrenformer and Voronoi RF suggests that training only on composition and coarse-grained structural features (spacegroup and Wyckoff positions in the case of Wrenformer; coordination numbers, local environment properties, etc. in the case of Voronoi RF) alone is insufficient to learn an extrapolatable map of the PES.
Given its strong performance on batch 1, it is possible that given sufficiently diverse training data, Wrenformer could become similarly accurate to the UIPs across the whole PES landscape at substantially less training and inference cost. However, the loss of predictive forces and stresses may make Wrenformer unattractive for certain applications even then.
A.8 Largest Errors vs. DFT Hull Distance
Given the large variety of models tested, we asked whether any additional insight into the errors can be gained from looking at how the predictions vary between different models. In fig. 9 we see two distinct groupings emerge when looking at the 200 structures with the largest errors. This clustering is particularly apparent when points are colored by model disagreement.
A.9 MEGNet formation energies from UIP-relaxed structures
Interestingly, table 4 shows that MEGNet RS2RE performs worse than CHGNet and M3GNet even when it has access to the ground truth DFT relaxed structures when making relaxed energy predictions. While the regression and classification metrics for MEGNet RS2RE are better than M3GNet→MEGNet and CHGNet→MEGNet, the difference is surprisingly small, especially compared to the regression performance of standalone MEGNet. This suggests that for downstream use, at least by an imprecise method such as another ML model, the UIP-relaxed structures are almost as helpful as the DFT-relaxed ones.
CHGNet, M3GNet and MEGNet were all trained on slightly different targets. While MEGNet was trained to predict formation energies, M3GNet and CHGNet were trained on raw DFT energies with the important distinction that CHGNet targets include MP2020 energy correction scheme [52] while M3GNet targets do not.
Changing the target from raw DFT energies to formation energies in principle merely constitutes a change of gauge since the difference is just a linear transformation of subtracting the elemental reference energies weighted by composition. But in practice, one might expect the problem of stability prediction to be easier when trained on formation energies, as the model output is one less step removed from the final target of convex hull distance. However, empirical evidence so far suggests the opposite. Both UIP→MEGNet combos perform worse than those same UIPs by themselves. There are too many confounding effects at play to draw firm conclusions but this tentatively suggests using formation energy as targets offers no benefits over raw DFT.
A.10 Exploratory Data Analysis
To give high-level insights into the MP training and WBM test set used in this work, we include element distributions for structures in both data sets (fig. 11). To show how frame selection from MP structure relaxation affected relative elemental abundance between MP relaxed structures and the snapshots in MPtrj, fig. 12 shows element occurrence ratios between MPtrj and MP. Believing MPtrj to be an influential dataset for the near-term continued development of universal interatomic potentials, we plot histograms showing the distributions of target values for energies, forces, stresses and magnetic moments in fig. 14. Similarly, fig. 15 plots the distribution of formation energies and convex hull distances (with respect to the convex hull spanned by MP materials only) of the WBM test set. This plot not only gives insight into the nature of the dataset but also emphasizes the increased difficulty of stability vs formation energy prediction arising from the much narrower distribution of convex hull distances compared to the more spread-out formation energy distribution.
To get a visual sense of coverage of atomic configuration space afforded by WBM compared to MP (as an indicator for the degree of out-of-distribution extrapolation required), we project matminer [53] features for all initial structures as well as relaxed MP structures onto two UMAP dimensions in fig. 10.
Finally, fig. 12(a) shows the elements-per-structure distribution of MP, MPtrj and WBM, normalized by dataset size. The mode of all three datasets is 3, but WBM’s share of ternary phases is noticeably more peaked than MP’s, which includes small numbers of unary and senary phases. The histogram in fig. 12(b) shows the distribution of the number of sites in MPtrj structures. The inset displays the same histogram log-scaled y-axis as well as a cumulative line to show that 90% of MPTrj structures contain fewer than 70 sites.
A.11 Predicted Formation Energy Analysis
We show the distribution for all materials in the WBM test set of model-predicted formation energies vs DFT ground truth in fig. 17 and the distribution of convex hull distance errors projected onto elements by composition for MACE and CHGNet in fig. 18.

### v3

@@
 Abstract
-Matbench Discovery simulates using machine learning (ML) energy models as pre-filters to DFT in a high-throughput search for stable inorganic crystals.
-We address the disconnect between (i) thermodynamic stability and formation energy and (ii) in-domain vs out-of-distribution performance.
-Alongside this paper, we publish a Python package to aid with future model submissions and a growing online leaderboard with further insights into trade-offs between various performance metrics.
-To answer the question of which ML methodology performs best at materials discovery, our initial release includes random forests, graph neural networks (GNN), one-shot predictors, iterative Bayesian optimizers and universal interatomic potentials (UIP).
-Ranked best-to-worst by their test set F1 score on thermodynamic stability prediction, we find MACE > CHGNet > M3GNet > ALIGNN > MEGNet > CGCNN > CGCNN+P > Wrenformer > BOWSR > Voronoi fingerprints with random forest.
-The top 3 models are UIPs which we establish to be by far the winning methodology for ML-guided materials discovery, achieving F1 scores of 0.6 for crystal stability classification and discovery acceleration factors (DAF) of up to 5x on the first 10k most stable predictions compared to dummy selection.
-We also highlight a disconnect between commonly used regression metrics and more task-relevant classification metrics.
-Accurate regressors are susceptible to unexpectedly high false-positive rates if those accurate predictions lie close to the decision boundary at above the convex hull where most materials are.
-Finally, we share valuable insights for maintainers of high throughput materials databases by demonstrating that universal potentials have matured enough to triage and determine compute budget allocation in future expansion efforts.
+The rapid adoption of machine learning (ML) in domain sciences necessitates best practices and standardized benchmarking for performance evaluation.
+We present Matbench Discovery, an evaluation framework for ML energy models, applied as pre-filters for high-throughput searches of stable inorganic crystals.
+This framework addresses the disconnect between thermodynamic stability and formation energy, as well as retrospective vs.
+prospective benchmarking in materials discovery.
+We release a Python package to support model submissions and maintain an online leaderboard, offering insights into performance trade-offs.
+To identify the best-performing ML methodologies for materials discovery, we benchmarked various approaches, including random forests, graph neural networks (GNNs), one-shot predictors, iterative Bayesian optimizers, and universal interatomic potentials (UIP).
+Our initial results rank models by test set F1 scores for thermodynamic stability prediction: EquiformerV2 + DeNS > Orb > SevenNet > MACE > CHGNet > M3GNet > ALIGNN > MEGNet > CGCNN > CGCNN+P > Wrenformer > BOWSR > Voronoi fingerprint random forest.
+UIPs emerge as the top performers, achieving F1 scores of 0.57–0.82 and discovery acceleration factors (DAF) of up to 6x on the first 10k stable predictions compared to random selection.
+We also identify a misalignment between regression metrics and task-relevant classification metrics.
+Accurate regressors can yield high false-positive rates near the decision boundary at 0 eV/atom above the convex hull.
+Our results demonstrate UIPs’ ability to optimize computational budget allocation for expanding materials databases.
+However, their limitations remain underexplored in traditional benchmarks.
+We advocate for task-based evaluation frameworks, as implemented here, to address these limitations and advance ML-guided materials discovery.
 1 Department of Physics, University of Cambridge, UK
@@
 3 Federal Institute of Materials Research and Testing (BAM), Germany
-4 Molecular Foundry, Lawrence Berkeley National Laboratory, USA
 4 Department of Materials Science and Engineering, University of California - Berkeley, USA
 1 Introduction
-Material science can be viewed as a combinatorial problem of mixing and arranging different atoms to leverage the complex range of properties that emerge.
-To date, combinations have been tested experimentally [1, 2] and have been simulated [3, 4, 5, 6].
-[7] identified possible quaternary materials allowed by electronegativity and charge-balancing rules.
+The challenge of evaluating and benchmarking the rapid evolution of machine learning models is common to many domain sciences.
+Specifically, the lack of pre-agreed upon tasks and data sets can obscure the performance of the model, making comparisons difficult.
+Material science presents one such domain, where in the last decade, the number of machine learning publications and associated models have increased dramatically.
+Similar to other domains, such as drug discovery and protein design, the ultimate success is often associated with the discovery of a new material with specific functionality.
+In the combinatorial sense, material science can be viewed as an optimization problem of mixing and arranging different atoms with a merit function that captures the complex range of properties that emerge.
+To date, combinations have been tested experimentally [bergerhoff_inorganic_1983, belsky_new_2002] , have been simulated[jain_commentary_2013, saal_materials_2013, curtarolo_aflow_2012, draxl_nomad_2018, schmidt2024improving], and upwards of possible quaternary materials are allowed by electronegativity and charge-balancing rules [davies_computational_2016].
 The space of quinternaries and higher is even less explored, leaving vast numbers of potentially useful materials to be discovered.
@@
 Any speed-up new discovery methods might yield should be leveraged to their fullest extent.
-Despite significant advances in empirical, theoretical and computational materials science, discovering new materials still requires complex calculations, labor-intensive trial-and-error experimentation, and often happens fortuitously rather than through rational design.
-Machine learning (ML) methods efficiently extract and distill trends from huge datasets, and can handle high dimensionality, multiple objectives [8], uncertainty [9, 10, 11], and noisy or sparse data [12, 13], making them powerful additions to the computational materials science tool set.
-ML models are less accurate and reliable but orders of magnitude faster than ab initio simulation.
-This makes them most suitable for use in high-throughput (HT) searches to triage more expensive, higher-fidelity simulation methods.
-The use of neural networks for learning the Kohn-Sham density-functional theory (DFT) potential energy surface (PES) can be traced as far back as [14].
+Despite significant advances in empirical, theoretical, and computational materials science, discovering new materials still requires complex calculations, labor-intensive trial-and-error experimentation, and often happens fortuitously rather than through rational design.
+Machine learning (ML) methods efficiently extract and distill trends from large datasets, and can handle high dimensionality, multiple objectives [riebesell_pushing_2024], uncertainty [borg_quantifying_2022, goodall_rapid_2022, zhu_fast_2023], and noisy or sparse data [depeweg_decomposition_2017, bartel_critical_2020], making them powerful additions to augment the traditional computational materials science tool set.
+In particular, we focus on the role of ML to accelerate the use of Kohn-Sham density-functional theory (DFT) in the materials discovery pipeline.
+In comparison to other simulation frameworks DFT offers a compelling compromise between fidelity and cost that has seen it adopted as a workhorse method by the computational material science community.
+The great strengths of DFT as a methodology have led it to demand up to 45% of core hours at the UK-based Archer2 Tier 1 supercomputer [ukriGoldilocksConvergence] and over 70% allocation time in materials science sector at NERSC [griffin2022computational, NERSC:2018].
+This heavy resource requirement drives demand for ways to reduce or alleviate its computational burden, such as efficiency improvements or substitution from ML approaches.
+ML models are less accurate and reliable but orders of magnitude faster than ab-initio simulation.
+This makes them most suitable for use in high-throughput (HT) searches to triage more expensive, higher-fidelity simulation methods such as DFT.
+The use of neural networks for learning the DFT potential energy surface (PES) can be traced as far back as [behler_generalized_2007].
 This work kicked off rapid advances and significant efforts to fit ever more sophisticated ML models to known samples of the PES.
-Initially, most of these models were trained and deployed as interatomic potentials (also known as force fields) to study known materials of interest, a workflow that requires curating custom training data for each new system of interest [15, 16].
-As larger and more diverse datasets have emerged from initiatives like the Materials Project (MP) [3], AFLOW [5] or the Open Quantum Materials Database (OQMD) [4], researchers have begun to train so-called universal models that cover 90 or more of the most-application relevant elements in the periodic table.
+Initially, most of these models were trained and deployed as interatomic potentials (also known as force fields) to study known materials of interest, a workflow that requires curating custom training data for each new system of interest [bartok_machine_2018, deringer_general-purpose_2020].
+As larger and more diverse datasets have emerged from initiatives like the Materials Project (MP) [jain_commentary_2013], AFLOW [curtarolo_aflow_2012] or the Open Quantum Materials Database (OQMD) [saal_materials_2013], researchers have begun to train so-called universal models that cover 90 or more of the most-application relevant elements in the periodic table.
 This opens up the prospect of ML-guided materials discovery to increase the hit rate of stable crystals and speed up DFT- and expert-driven searches.
@@
 As ML models have grown in complexity and applicability, benchmark datasets need to grow with them to accurately measure their usefulness.
-However, due the the rapid pace of the field and the variety of possible approaches for framing the discovery problem, no large-scale benchmark yet exists for measuring the ability of ML to accelerate materials discovery.
+However, due to the rapid pace of the field and the variety of possible approaches for framing the discovery problem, no large-scale benchmark yet exists for measuring the ability of ML to accelerate materials discovery.
 As a result, it is unclear which methodologies or models are best suited for this task.
-Recent works focusing on prospective computational materials discovery have proposed strategies based on one-shot coordinate free predictors [10], iterative Bayesian optimizers [17], and universal interatomic potentials (UIP) [18, 19, 20].
-These papers deploy their respective model on specific systems and custom datasets to highlight certain strengths but have yet to be compared in a systematic standardized manner that allows them to be ranked by their ability to accelerate materials discovery.
+Recent works focusing on prospective computational materials discovery have proposed strategies based on one-shot coordinate free predictors [goodall_rapid_2022], iterative Bayesian optimizers [zuo_accelerating_2021], and universal interatomic potentials (UIP) [chen_universal_2022, deng_chgnet_2023, batatia_mace_2023].
+These papers deploy their respective models on specific systems and custom datasets to highlight certain strengths but have yet to be compared in a systematic standardized manner that allows them to be ranked by their ability to accelerate materials discovery.
 Our work aims to identify the state-of-the-art (SOTA) model by proposing a novel evaluation framework that closely simulates a real-world discovery campaign guided by ML models.
-Specifically, we designed a benchmark task that tackles three central challenges. We believe these challenges to be requirements when seeking to justify experimental validation of ML predictions:
+Our findings show that UIPs outperform all other methodologies we tested both in accuracy and robustness.
+We hope that creation of benchmarks following this framework create a pathway through which interdisciplinary researchers with limited material science backgrounds can contribute usefully to model architecture and methodology development on a relevant task and thereby aid progress in Material Science.
+2 Evaluation Framework for Materials Discovery
+In this work we proposed a benchmark task that addresses four central challenges that we believe to be requirements when seeking to justify experimental validation of ML predictions:
 1.
-Mimic Real Applications: Idealized and overly simplified benchmarks can fail to reflect the real-world challenges a model faces when used in an actual discovery campaign.
+Prospective Benchmarking: Idealized and overly simplified benchmarks can fail to reflect the real-world challenges a model faces when used in an actual discovery campaign.
 This can result in a disconnect between benchmark metrics and real-world performance.
-Possible reasons for this include choosing the wrong target [13] or picking an unrepresentative train/test split [21, 22].
-In the case of materials discovery, formation energies - although widely used as regression targets - do not indicate thermodynamic stability.
-Distance to the convex hull spanned by competing phases in the same chemical system is a much more informative indicator of thermodynamic stability.
+Possible reasons for this include choosing the wrong target [bartel_critical_2020] or picking an unrepresentative training and test splits [wu_moleculenet_2018, kpanou_robustness_2021].
+For small datasets of materials properties, ‘Leave-Out’ data splitting strategies are often used to assess model performance [meredig2018can, cubuk2019screening, zahrt2020cautionary].
+However, in our target domain large quantities of diverse data () are available and hence retrospective splitting strategies predicated on clustering can end up testing artificial or unrepresentative use cases.
+This encourages using new sources of prospectively generated test data to understand application performance.
+Adopting this principle, the intended discovery workflow should be used to generate the test data leading to a substantial but realistic covariate shift between the training and test distributions that gives a much better indicator likely of performance on additional application of the same discovery workflow.
+Relevant Targets: In the case of materials discovery, high-throughput DFT calculated formation energies - although widely used as regression targets - do not indicate thermodynamic stability or synthesizability.
+In contrast, the thermodynamic competition between one material and its competing phases in the same chemical system determines its thermodynamic stability, which can be quantified by the material’s energy to the convex hull phase diagram.
+The distance to the convex hull is the leading order term and strongly predictive of (meta-)stability at standard conditions [sun2016thermodynamic].
+Whilst other factors such as finite temperature effects, vibrational, and configurational entropy can modify a material’s stability under real-world conditions they are prohibitive to calculate in high-throughput workflows, leaving the distance to the convex hull as the best signal-vs-data-availability compromise to use as a target.
 Moreover, ML models relying on relaxed crystal structures as input render any discovery pipeline circular since obtaining relaxed structures requires computationally expensive DFT simulations, thereby depending on the very process we intend to accelerate.
 2.
-Opportunity cost: Accurate regressors are susceptible to unexpectedly high false-positive rates if those accurate predictions lie close to the decision boundary.
-Looking purely at global metrics like MAE, RMSE and can give practitioners a false sense of security about their model’s reliability.
-Failed experiments incur a high opportunity cost by wasting lab time and resources.
+Informative Metrics: Global error metrics like MAE, RMSE and can often give practitioners a false sense of security about the model’s reliability.
+For example, accurate regressors are susceptible to unexpectedly high false-positive rates if those nominally accurate predictions lie close to the decision boundary and the failed predictions can incur a high opportunity cost by wasting lab time and resources.
+Consequently, models should be measured on their ability to lead to correct decision making patterns not regression accuracy.
+One way to do this is by defining a selection criteria and then evaluating regression models primarily in terms of their classification performance.
 3.
-Scalability: Future discovery efforts are likely to encounter large data regimes.
-Small benchmarks can lack chemical diversity, obfuscate poor scaling relations or poor out-of-distribution (OOD) performance (see fig.
-10 for a UMAP projection of train and test sets).
-For instance, random forests achieve excellent performance on small datasets but are typically outperformed by neural networks on large datasets due to the benefits of representation learning [23].
-Two prior efforts worth highlighting that have partially addressed the above challenges are Matbench [24] and the Open Catalyst Project (OCP) [25].
+Scalability: Future materials discovery efforts are likely to target broad chemical spaces and large data regimes.
+Small benchmarks can lack chemical diversity, and obfuscate poor scaling relations or weak out-of-distribution (OOD) performance.
+For instance, random forests achieve excellent performance on small datasets but are typically outperformed by neural networks on large datasets due to the benefits of representation learning [goodall_predicting_2020].
+Whilst we propose that large training sets are necessary to adequately differentiate the ability of models to learn in the larger data regime, given the enormous size of the configurational space of materials yet to be explored we also propose that it is important to construct a task where the test set is larger than the training set to mimic true deployment at scale.
+No other inorganic materials benchmarks test the prospects of large-scale deployment in this manner.
+We highlight two specific benchmarking efforts that have partially addressed the above challenges: Matbench [dunn_benchmarking_2020] and the Open Catalyst Project (OCP) [chanussot_open_2021].
+Other valuable efforts such as MatSciML [lee2023matsciml] and JARVIS-Leaderboard [choudhary2024jarvis] aggregate a wide variety of material science related benchmark tasks, including from Matbench and OCP, but do not introduce distinct benchmarking design patterns to those seen in Matbench or the OCP.
 By providing a standardized collection of 13 datasets ranging in size from 300 to 132,000 samples from both DFT and experimental sources, Matbench addresses the scalability challenge, highlighting how model performance changes as a function of data regime.
@@
 We believe a task that simulates a materials discovery campaign by requiring materials stability prediction from unrelaxed structures to be a missing piece here.
-OCP is a large-scale initiative to discover substrate-adsorbate combinations that catalyze key industrial reactions that process said adsorbates into more useful products.
-The OCP has released two data sets thus far, OCP20 [25] and OCP22 [26], for training and benchmarking ML models.
+OCP is a large-scale initiative aimed at discovering substrate-adsorbate combinations that can catalyze critical industrial reactions, transforming these adsorbates into more useful products.
+The OCP has released two datasets thus far, OCP20 [chanussot_open_2021] and OCP22 [tran_open_2022] for training and benchmarking ML models.
 OCP certainly addressed challenge 1 of closely mimicking a real-world problem.
-They have recently shown that despite not reaching their target accuracy to entirely replace DFT, using ML in conjunction with confirmatory DFT calculations dramatically speeds up their combinatorial screening workflow [27].
-We believe that addressing these three challenges will enable future ML-guided discovery efforts to expand materials databases to confidently select appropriate models and methodology.
-Our findings show that UIPs outperform all other methodologies we tested both in accuracy and robustness.
-1.1 Evaluation Framework for Materials Discovery
-We propose a novel evaluation framework that places no constraints on the type of data a model trains on as long as it would be available to a practitioner conducting a real materials discovery campaign.
+They have recently shown that despite not reaching their target accuracy to entirely replace DFT, using ML in conjunction with confirmatory DFT calculations dramatically speeds up their combinatorial screening workflow [lan_adsorbml_2023].
+The team behind the OCP has a second initiative targeting materials for direct air capture called OpenDAC that has shared the ODAC23 dataset [sriram2024open].
+The OpenDAC benchmark is setup identically to the OCP.
+We believe that addressing these four challenges will result in benchmarks that enable future ML-guided discovery efforts to confidently select appropriate models and methodologies for the expansion of computational materials databases.
+2.1 Matbench Discovery
+We propose a novel evaluation framework that places no constraints on the type of data a model is trained on as long as it would be available to a practitioner conducting a real materials discovery campaign.
+This means that for the high-throughput DFT data considered, any subset of the energies, forces, stresses or any other properties that can be routinely extracted from DFT calculations, such as magnetic moments, are valid training targets.
+All of these would be available to a practitioner performing a real materials discovery campaign and hence are permitted for training any model submission.
 We only enforce that at test time, all models must make predictions on the convex hull distance of the relaxed structure with only the unrelaxed structure as input.
 This setup avoids circularity in the discovery process, as unrelaxed structures can be cheaply enumerated through elemental substitution methodologies and do not contain information inaccessible in a prospective discovery campaign.
-We choose to focus on the relaxed structure’s convex hull distance as a measure of thermodynamic stability rather than the formation energy as it informs the decision on whether to pursue a potential candidate crystal.
-This decision was also motivated by [13] who found even composition-only models capable of predicting DFT formation energies with useful accuracy.
+Figure 1 provides a visual overview of design choices.
+The convex hull distance of a relaxed structure is chosen as the measure of its thermodynamic stability, rather than the formation energy, as it informs the decision on whether to pursue a potential candidate crystal.
+This decision was also motivated by [bartel_critical_2020] who found that even composition-only models are capable of predicting DFT formation energies with useful accuracy.
 However, when tasking those same models with predicting decomposition enthalpy, performance deteriorated sharply.
-This insight meant that ML models are much less useful than DFT for discovering new inorganic solids.
-Moreover, the qualitative leap in performance from Roost [23], the best compositional model benchmarked, to CGCNN [28], the single structural model they tested, highlights that structure plays a crucial role in determining material stability.
-Hence our decision to restrict test-time model input to unrelaxed structures to avoid the above-mentioned circularity and measure true prospective utility while still allowing for subtle atomic-configuration-informed energy estimates.
-We recognize that our ground truth is ignorant of entropic stabilization and metastable states.
-While these factors influence experimental synthesizability, we have to neglect them when simulating an HT search since the convex hull distance as predicted by zero-temperature DFT is the best proxy of true crystal stability available for large datasets.
+This insight exposes how ML models are much less useful than DFT for discovering new inorganic solids than would be expected given their low prediction errors for formation energies due to the impact of random as opposed to systematic errors.
 Standard practice in ML benchmarks is to hold all variables fixed – most importantly the training data – and vary only the model architecture to isolate architectural effects on the performance.
@@
 This allows them to learn a fundamentally higher fidelity model of the physical interactions between ions.
-That is a genuine advantage of the architecture and something any benchmark aiming to identify the optimal methodology for materials discovery must reflect, as ours does.
-As a result, our benchmark contains models trained on varying data sets.
-We define the Materials Project (MP) [3] database release (v2022.10.28 as the maximum allowed training set for any valid model submission.
-Any subset of the energies, forces, stresses, magnetic moments or any other recorded DFT data are valid training targets.
-All of these would be available to a practitioner performing a real materials discovery campaign and hence are permitted for any model submission.
-Models may train on the complete set of relaxation frames, or any subset thereof such as the final relaxed structures only.
-Our test set consists of the unrelaxed structures in the WBM dataset [29].
+That is a genuine advantage of the architecture and something any benchmark aiming to identify the optimal methodology for materials discovery must reflect.
+In light of this utilitarian perspective, our benchmark contains models trained on varying datasets, and any model that can intake more physical modalities from DFT calculations is a valid model for materials discovery.
+We define the Materials Project (MP) [jain_commentary_2013] v2022.10.28 database release as the maximum allowed training set for any compliant model submission.
+Models may train on the complete set of relaxation frames, or any subset thereof such as the final relaxed structures.
+Any subset of the energies, forces, and stresses are valid training targets.
+In addition any auxiliary tasks such as predicting electron densities, magnetic moments, site-partitioned charges, etc.
+that can be extracted from the output of the DFT calculations are allowed for multi-task learning [shoghi2023molecules].
+Our test set consists of the unrelaxed structures in the WBM dataset [wang_predicting_2021].
 Their target values are the PBE formation energies of the corresponding DFT-relaxed structures.
 2.1.1 Materials Project Training Set
-The Materials Project is a well-known effort to calculate the properties of all inorganic materials using high-throughput ab initio methods.
-Seeded from a subset of the Inorganic Crystal Structure Database (ICSD) [30], the initial release of the database consisted of 9 k crystals.
-At the time of writing, the Materials Project database [3] has grown to 154 k crystals, covering diverse chemistries and providing relaxed and initial structures as well as the relaxation trajectory for every entry.
+The Materials Project is a widely-used database of inorganic materials properties that have been calculated using high-throughput ab-initio methods.
+At the time of writing, the Materials Project database [jain_commentary_2013] has grown to 154 k crystals, covering diverse chemistries and providing relaxed and initial structures as well as the relaxation trajectory for every entry.
 Our benchmark defines the training set as all data available from the v2022.10.28 MP release.
 We recorded a snapshot of energies, forces, stresses and magnetic moments for all MP ionic steps on 2023-03-15 as the canonical training set for Matbench Discovery, and provide convenience functions through our Python package for easily feeding that data into future model submissions to our benchmark.
-The flexibility in specifying the maximum allowed data set is intended to allow authors to experiment with and exploit the large variety of available data.
+Flexibility in specifying the dataset allows authors to experiment with and fully exploit the available data.
 This choice is motivated by two factors.
-First, whilst it may appear that models that train on multiple snapshots containing energies, forces and stresses are getting more training data than models that are only trained or able to be trained on the energies of the relaxed structures the critical factor is that all this additional data was generated as a byproduct of the workflow to produce the relaxed structures i.e.
-all models have been trained from the same total cost of training data acquisition.
+First, it may seem that models trained on multiple snapshots containing energies, forces, and stresses receive more training data than models trained only on the energies of relaxed structures.
+However, the critical factor is that all this additional data was generated as a byproduct of the workflow to produce relaxed structures.
+Consequently, all models are being trained using data acquired at the same overall cost.
 If some architectures or approaches can leverage more of this byproduct data to make improved predictions this is a fair comparison between the two models.
@@
 composition-only vs structure-available in Matbench or IS2RS, IS2RE, S2EF in OCP) and do not make direct comparisons of this manner.
-Second, recent work in the space from [31, 32] has claimed that much of the data in large databases like MP are redundant and that models can be trained more effectively by taking a subset of these large data pools.
-We believe that from a systems-level perspective identification of novel cleaning and or active-learning strategies that do not leak signal from the test set are equally as important as architectural improvements as they can lead to similar improvements in performance, particularly given the prevalence of errors in high-throughput DFT that can disrupt learning of the PES.
+Second, recent work in the space from [li_critical_2023, li_exploiting_2023] has claimed that much of the data in large databases like MP are redundant and that models can be trained more effectively by taking a subset of these large data pools.
+From a systems-level perspective, identifying novel cleaning or active-learning strategies to make better use of available data may be as crucial as architectural improvements, as both can similarly enhance performance, especially given the prevalence of errors in high-throughput DFT.
 Consequently, such strategies where they lead to improved performance should be able to be recognized within the benchmark.
 We encourage submissions to submit ablation studies showing how different system-level choices affect performance.
-We highlight for example data sets that are valid within the rules of the benchmark that take advantage of these freedoms.
-The first is the MPF.2021.2.8 data set [18] curated to train the M3GNet model which takes a subset of materials from the v2021.02.08 MP release.
+Another example of a system-level choice that may impact performance is the choice of optimizer, for example FIRE [bitzek2006structural] vs L-BFGS, in the relaxation when using UIP models.
+We highlight several example datasets that are valid within the rules of the benchmark that take advantage of these freedoms.
+The first is the MP-crystals-2019.4.1 dataset [chen_graph_2019] which is a subset of 133,420 crystals and their formation energies that form a subset of the v2021.02.08 MP release.
+The MP-crystals-2022.10.28 dataset is introduced with this work comprising a set of 154,719 structures and their formation energies drawn from the v2021.02.08 MP release.
+The next is the MPF.2021.2.8 dataset [chen_universal_2022] curated to train the M3GNet model which takes a subset of 62,783 materials from the v2021.02.08 MP release.
+The curators of the MPF.2021.2.8 dataset down-sampled the v2021.02.08 release significantly to select a subset of calculations that they believed to be most self-consistent.
 Rather than taking every ionic step from the relaxation trajectory this dataset opts to select only the initial, final and one intermediate structure for each material to avoid biasing the dataset towards examples where more ionic steps were needed to relax the structure.
-Additionally, the curators of the MPF.2021.2.8 data set down-sampled the v2021.02.08 release significantly to select a subset of calculations that they believed to be most self-consistent.
+Consequently the dataset consists of 188,349 structures.
 The MPF.2021.2.8 is a proper subset of the training data as no materials were deprecated between the v2021.02.08 and v2022.10.28 database releases.
-The second data set we highlight, with which several of the UIP models have been trained, is the MPtrj dataset [19].
+The final dataset we highlight, with which several of the UIP models have been trained, is the MPtrj dataset [deng_chgnet_2023].
 This dataset was curated from the earlier v2021.11.10 MP release.
-The MPtrj dataset is a proper subset of the allowed training data but several potentially anomalous examples from within MP were cleaned out of the data set before the frames were subsampled to remove.
+The MPtrj dataset is a proper subset of the allowed training data but several potentially anomalous examples from within MP were cleaned out of the dataset before the frames were subsampled to remove redundant frames.
 It is worth noting that the v2022.10.28 release contains a small number of additional Perovskite structures not found in MPtrj that could be added to the training set within the scope of the benchmark.
@@
 2.1.2 WBM Test Set
-The WBM data set [29] consists of 257 k structures generated via chemical similarity-based elemental substitution of MP source structures followed by DFT relaxation and calculating each crystal’s convex hull distance.
-Which element replaces an existing one in a given source structure was determined by random sampling according to the weights in a chemical similarity matrix data-mined from the ICSD [33].
-The WBM authors performed 5 iterations of this substitution process (we refer to these as batches).
+The WBM dataset [wang_predicting_2021] consists of 257,487 structures generated via chemical similarity-based elemental substitution of MP source structures followed by DFT relaxation and calculating each crystal’s convex hull distance.
+The element substitutions applied to a given source structure were determined by random sampling according to the weights in a chemical similarity matrix data-mined from the ICSD [glawe_optimal_2016].
+The WBM authors performed 5 iterations of this substitution process (we refer to these steps as batches).
 After each step, the newly generated structures found to be thermodynamically stable after DFT relaxation flow back into the source pool to partake in the next round of substitution.
-This split of the data into batches of increasing substitution count is a unique and compelling feature of the test set as it allows out-of-distribution (OOD) testing by seeing if model performance degrades with substitution count.
-A higher number of elemental substitutions on average carries the structure further away from the region of material space covered by the MP training set (see figs.
-8 and 10 for details).
-Whilst this batch information makes the WBM data set an exceptionally useful data source for examining the extrapolation performance of ML models, our evaluation primarily looks at metrics that consider all batches as a single test set.
+This split of the data into batches of increasing substitution count is a unique and compelling feature of the test set as it allows out-of-distribution (OOD) testing by examining whether model performance degrades for later batches.
+A higher number of elemental substitutions on average carries the structure further away from the region of material space covered by the MP training set (see fig.
+S6 for details).
+Whilst this batch information makes the WBM dataset an exceptionally useful data source for examining the extrapolation performance of ML models, we look primarily at metrics that consider all batches as a single test set.
+In order to control for the potential adverse effects of leakage between the MP training set and the WBM test set, we cleaned the WBM test set based on protostructure matching.
+We refer to the combination of a materials prototype and the elemental assignment of its wyckoff positions as a protostructure following [parackal2024identifying].
+First we removed 524 pathological structures in WBM based on formation energies being larger than 5 eV/atom or smaller than -5 eV/atom.
+We then removed from the WBM test set all examples where the final protostructure of a WBM material matched the final protostructure of an MP material.
+In total 11,175 materials were cleaned using this filter.
+We further removed all duplicated protostructures within WBM, keeping the lowest energy structure in each instance, leaving 215,488 structures in the unique prototype test set.
 Throughout this work, we define stability as being on or below the convex hull of the MP training set ().
-42k out of 257k materials in WBM satisfy this criterion.
+32,942 out of 215,488 materials in WBM unique prototype test set satisfy this criterion.
 Of these, 33k are unique prototypes, meaning they have no matching structure prototype in MP nor another higher-energy duplicate prototype in WBM.
 Our code treats the stability threshold as a dynamic parameter, allowing for future model comparisons at different thresholds.
-This may reveal systematic differences in stability calibration, as hinted by the different TPR/TNR tradeoffs observed for CHGNet and M3GNet in table 3.
-A looser stability threshold (e.g.
-) might benefit a model like M3GNet which is more conservative in its stability predictions, whereas a stricter threshold may benefit a model like CHGNet that is more optimistic in its stability predictions.
 For initial analysis in this direction, see fig.
-3 in the SI.
+S1 in the SI.
 As WBM explores regions of materials space not well sampled by MP, many of the discovered materials that lie below MP’s convex hull are not stable relative to each other.
-Of the 42k that lie below the MP convex hull less than half, or around 20k, remain on the joint MP+WBM convex hull.
+Of the 33k that lie below the MP convex hull less than half, or around 20k, remain on the joint MP+WBM convex hull.
 This observation suggests that many WBM structures are repeated samples in the same new chemical spaces.
-It also highlights a critical aspect of this benchmark in that we purposely operate on an incomplete convex hull.
+It also highlights a critical aspect of this benchmark in that we knowingly operate on an incomplete convex hull.
 Only current knowledge of competing points on the PES is accessible to a real discovery campaign and our metrics are designed to reflect this.
+2.1.3 Limitations of this Framework
+Whilst the framework proposed here directly mimics a common computational materials discovery workflow, it is worth highlighting that there still exist significant limitations to these traditional computational workflows that can prevent the material candidates suggested by such a workflow from being able to be synthesized in practice.
+For example, high-throughput DFT calculations often use small unit cells which can lead to artificial orderings of atoms.
+The corresponding real material may be disordered due to entropic effects that cannot be captured in the zero-kelvin thermodynamic convex hull approximated by DFT [cheetham2024artificial].
+Another issue is that, when considering small unit cells, the DFT relaxations may get trapped at dynamically unstable saddle points in the true PES.
+This failure can be detected by calculating the phonon spectra for materials predicted to be stable.
+However, the cost of doing so with DFT is often deemed prohibitive for high-throughput screening.
+The lack of information about the dynamic stability of nominally stable materials in the WBM test set prevents this work from considering this important criteria as an additional screening filter.
+However, recent progress in the development of UIPs suggests that ML approaches will soon provide sufficiently cheap approximations of these terms for high-throughput searches [batatia2023foundation, deng2024overcoming].
+As the task presented here begins to saturate, we believe that future discovery benchmarks should extend upon the framework proposed here to also incorporate criteria based on dynamic stability.
+[…]
+We design the benchmark considering a positive label for classification as being on or below the convex hull of the MP training set.
+An alternative formulation would be to say that materials in WBM that are below the MP convex hull but do not sit on the combined MP+WBM convex hull are negatives.
+The issue with such a design is that it involves unstable evaluation metrics.
+If we consider the performance against the final combined convex hull rather than the initial MP convex hull, then each additional sample considered can retroactively change whether or not a previous candidate would be labeled as a success as it may no-longer sit on the hull.
+Since constructing the convex hull is computationally expensive, this path dependence makes it impractical to evaluate cumulative precision metrics (see Figure 2).
+The chosen setup does increase the number of positive labels and could consequently be interpreted as overestimating the performance.
+This overestimation decreases as the convex hull becomes better sampled.
+Future benchmarks building on this work could make use of combination of MP+WBM to control this artifact.
+An alternative framework could report metrics for each WBM batch in turn and retrain between batches, this approach was undesirable here as it increases the cost of submission five-fold and introduces many complexities, for example, should each model only retrain on candidates it believed to be positive, that would make fair comparison harder.
 2.2 Models
-To test a wide variety of methodologies proposed for learning the potential energy landscape, our initial benchmark release includes 10 models.
+To test a wide variety of methodologies proposed for learning the potential energy landscape, our initial benchmark release includes 13 models.
+Next to each model’s name we give the training targets that were used: E - Energy, F - Forces, S - Stresses and M - Magnetic moments.
+The subscripts G and D refer to whether gradient-based or direct prediction methods were used to obtain force and stress predictions.
 1.
-MACE [20] (UIP-GNN) - MACE builds upon the recent advances [34, 35] in equivariant neural network architectures by proposing an approach to computing high-body-order features efficiently via Atomic Cluster Expansion [36].
+EquiformerV2 + DeNS [liao2024equiformerV2, liao2024generalizing] (EFSD) - EquiformerV2 builds on the first Equiformer model [liao2023equiformer] by replacing the convolutions with eSCN convolutions [passaro2023escn] as well as a range of additional tweaks to make better use of the ability to scale to higher using eSCN convolutions.
+EquiformerV2 uses direct force prediction rather than taking the forces as the derivative of the energy predictions for computational efficiency.
+Here we take the pre-trained "eqV2 S DeNS" [barroso2024open] trained on the MPTrj dataset.
+This model in addition to supervised training on energies, forces, and stresses makes use of a auxilary training task based on de-noising non-equilibrium structures [liao2024generalizing].
+We refer to this model as "EquiformerV2 + DeNS" in the text and "eqV2 S DeNS" in plots.
+2.
+Orb [neumann2024orb] (EFSD) - Orb is a lightweight model architecture developed to scale well for the simulation of large systems such as metal organic frameworks.
+Rather than constructing an architecture that is equivariant by default, Orb instead makes use of data augmentation during training to achieve approximate equivariance.
+This simplifies the architecture allowing for faster inference.
+We report results for the "orb-mptrj-only-v2" model which was pre-trained using a diffusion-based task on MPTrj before supervised training on the energies, forces and stresses in MPTrj.
+For simplicity we refer to this model as "ORB MPTrj".
+3.
+SevenNet [park2024scalable] (EFSG) - SevenNet emerged from an effort to improve the performance of message passing neural networks [gilmer2017neural] when used to conduct large scale simulations involving that benefit from parallelism via spatial decomposition.
+Here we use the pre-trained "SevenNet-0_11July2024" trained on the MPTrj dataset.
+The SevenNet-0 model is an equivariant architecture based on a NequIP [batzner_equivariant_2022] architecture that mostly adopts the GNoME [merchant2023scaling] hyper-parameters.
+SevenNet-0 differs from NequIP and GNoME by replacing the tensor product in the self-connection layer with a linear layer applied directly to the node features, this reduces the number of parameters from 16.24 million in GNoME to 0.84 million for SevenNet-0.
+For simplicity we refer to this model as "SevenNet".
+4.
+MACE [batatia_mace_2023] (EFSG) - MACE builds upon the recent advances [thomas_tensor_2018, batzner_equivariant_2022] in equivariant neural network architectures by proposing an approach to computing high-body-order features efficiently via Atomic Cluster Expansion [drautz_atomic_2019].
 Unlike the other UIP models considered MACE was primarily developed for molecular dynamics of single material systems and not the universal use case studied here.
-It is the only -equivariant model we tested.
-2.
-CHGNet [19] (UIP-GNN) - CHGNet is a UIP for charge-informed atomistic modeling.
-Its distinguishing feature is that it was trained to predict magnetic moments on top of energies, forces and stresses in the MPTrj dataset consisting of relaxation trajectories for 1.5 million MP structures (see section A.10 for detailed analysis of this dataset).
+The authors trained MACE on the MPTrj dataset, these models have been shared under the name "MACE-MP-0" [batatia2023foundation] and we report results for the "2023-12-03" version commonly called "MACE-MP-0 (medium)".
+For simplicity we refer to this model as "MACE".
+5.
+CHGNet [deng_chgnet_2023] (EFSGM) - CHGNet is a UIP for charge-informed atomistic modeling.
+Its distinguishing feature is that it was trained to predict magnetic moments on top of energies, forces and stresses in the MPtrj dataset (which was prepared for the purposes of training CHGNet).
 By modeling magnetic moments, CHGNet learns to accurately represent the orbital occupancy of electrons which allows it to predict both atomic and electronic degrees of freedom.
-3.
-M3GNet [18] (UIP-GNN) - M3GNet is a GNN-based universal interatomic potential (UIP) (as in covering the full periodic table) for materials trained on up to 3-body interactions in the initial, middle and final frame of MP DFT relaxations.
+We make use of the pre-trained "v0.3.0" CHGNet model from [deng_chgnet_2023].
+6.
+M3GNet [chen_universal_2022] (EFSG) - M3GNet is a GNN-based UIP for materials trained on up to 3-body interactions in the initial, middle and final frame of MP DFT relaxations.
 The model takes the unrelaxed input and emulates structure relaxation before predicting energy for the pseudo-relaxed structure.
-4.
-ALIGNN [37] (GNN) - The Atomistic Line Graph Neural Network (ALIGNN) is a message passing GNN architecture that takes as input both the interatomic bond graph and a line graph corresponding to 3-body bond angles.
+We make use of the pre-trained "v2022.9.20" M3GNet model from [chen_universal_2022] trained on the compliant MPF.2021.2.8 dataset.
+7.
+ALIGNN [choudhary_atomistic_2021] (E) - The Atomistic Line Graph Neural Network (ALIGNN) is a message passing GNN architecture that takes as input both the interatomic bond graph and a line graph corresponding to 3-body bond angles.
 The ALIGNN architecture involves a global pooling operation which means that it is ill-suited to force-field applications.
-To address this the ALIGNN-FF model was later introduced without global pooling [38].
-5.
-MEGNet [39] (GNN) - MatErials Graph Network is another GNN similar to CGCNN for material properties of relaxed structures that also updates the edge and global features (like pressure, temperature, entropy) in its message passing operation.
+To address this the ALIGNN-FF model was later introduced without global pooling [choudhary_unified_2023].
+We trained ALIGNN on the MP-crystals-2022.10.28 dataset for this benchmark.
+8.
+MEGNet [chen_graph_2019] (E) - MatErials Graph Network is another GNN-based architecture that also updates a set of edge and global features (like pressure and temperature) in its message passing operation.
 This work showed that learned element embeddings encode periodic chemical trends and can be transfer-learned from large datasets (formation energies) to predictions on small data properties (band gaps, elastic moduli).
-6.
-CGCNN [28] (GNN) - The Crystal Graph Convolutional Neural Network (CGCNN) was the first neural network model to directly learn 8 different DFT-computed material properties from a graph representing the atoms and bonds in a periodic crystal.
+We make use of the pre-trained "Eform_MP_2019" MEGNet model trained on the compliant MP-crystals-2019.4.1 dataset.
+9.
+CGCNN [xie_crystal_2018] (E) - The Crystal Graph Convolutional Neural Network (CGCNN) was the first neural network model to directly learn 8 different DFT-computed material properties from a graph representing the atoms and bonds in a periodic crystal.
 CGCNN was among the first to show that just like in other areas of ML, given large enough training sets, neural networks can learn embeddings that outperform human-engineered structure features directly from the data.
-7.
-CGCNN+P [40] (GNN) - This work proposes simple, physically motivated structure perturbations to augment stock CGCNN’s training data of relaxed structures with structures resembling unrelaxed ones but mapped to the same DFT final energy.
+We trained an ensemble of 10 CGCNN models on the MP-crystals-2022.10.28 dataset for this benchmark.
+10.
+CGCNN+P [gibson_data-augmentation_2022] (E) - This work proposes simple, physically motivated structure perturbations to augment stock CGCNN’s training data of relaxed structures with structures resembling unrelaxed ones but mapped to the same DFT final energy.
 Here we chose , meaning the training set is augmented with 5 random perturbations of each relaxed MP structure mapped to the same target energy.
@@
 The authors confirm this by demonstrating a lowering of the energy error on unrelaxed structures.
--
-8.
-Wrenformer (Transformer) - For this benchmark, we introduce Wrenformer which is a variation on the coordinate-free Wren model [10] constructed using standard QKV-self-attention blocks [41] in place of message-passing layers.
+We trained an ensemble of 10 CGCNN+P models on the MP-crystals-2022.10.28 dataset for this benchmark.
+11.
+Wrenformer (E) - For this benchmark, we introduce Wrenformer which is a variation on the coordinate-free Wren model [goodall_rapid_2022] constructed using standard QKV-self-attention blocks [vaswani_attention_2017] in place of message-passing layers.
 This architectural adaptation reduces the memory usage allowing the architecture to scale to structures with greater than 16 Wyckoff positions.
-Like its predecessor, Wrenformer is a fast coordinate-free model aimed at accelerating screening campaigns where even the unrelaxed structure is a priori unknown.
+Like its predecessor, Wrenformer is a fast coordinate-free model aimed at accelerating screening campaigns where even the unrelaxed structure is a priori unknown [parackal2024identifying].
 The key idea is that by training on the coordinate anonymized Wyckoff positions (symmetry-related positions in the crystal structure), the model learns to distinguish polymorphs while maintaining discrete and computationally enumerable inputs.
 The central methodological benefit of an enumerable input is that it allows users to predict the energy of all possible combinations of spacegroup and Wyckoff positions for a given composition and maximum unit cell size.
-The lowest-ranked prototypes can then be fed into downstream analysis or modeling.
--
-9.
-BOWSR [17] (BO-GNN) - BOWSR combines a symmetry-constrained Bayesian optimizer (BO) with a surrogate energy model to perform an iterative exploration-exploitation-based search of the potential energy landscape.
-Here we use MEGNet [39] for the energy model as proposed in the original work.
+The lowest-ranked protostructures can then be fed into downstream analysis or modeling.
+We trained an ensemble of 10 Wrenformer models on the MP-crystals-2022.10.28 dataset for this benchmark.
+12.
+BOWSR [zuo_accelerating_2021] (E) - BOWSR combines a symmetry-constrained Bayesian optimizer (BO) with a surrogate energy model to perform an iterative exploration-exploitation-based search of the potential energy landscape.
+Here we use the pre-trained "Eform_MP_2019" MEGNet model [chen_graph_2019] for the energy model as proposed in the original work.
 The high sample count needed to explore the PES with BO makes this by far the most expensive model tested.
-10.
-Voronoi RF [42] (Fingerprint) - A random forest trained to map a combination of composition-based Magpie features [43] and structure-based relaxation-robust Voronoi tessellation features (effective coordination numbers, structural heterogeneity, local environment properties, …) to DFT formation energies.
-This fingerprint-based model predates most deep learning for materials but significantly improved over earlier fingerprint-based methods such as the Coulomb matrix [44] and partial radial distribution function features [45].
+13.
+Voronoi RF [ward_including_2017] (E) - A random forest trained to map a combination of composition-based Magpie features [ward_general-purpose_2016] and structure-based relaxation-robust Voronoi tessellation features (effective coordination numbers, structural heterogeneity, local environment properties, …) to DFT formation energies.
+This fingerprint-based model predates most deep learning for materials but significantly improved over earlier fingerprint-based methods such as the Coulomb matrix [rupp_fast_2012] and partial radial distribution function features [schutt_how_2014].
 It serves as a baseline model to see how much value the learned featurization of deep learning models can extract from the increasingly large corpus of available training data.
+We trained Voronoi RF on the MP-crystals-2022.10.28 dataset for this benchmark.
 3 Results
-Table 1 shows performance metrics for all models included in the initial release of Matbench Discovery.
-MACE takes the top spot and emerges as the current SOTA for ML-guided materials discovery.
-It outperforms the 9 other models on all 6 classification metrics but is slightly overtaken by CHGNet in 2 of 3 regression metrics, and RMSE, which are sensitive to outliers.
-The discovery acceleration factor (DAF) measures how many more stable structures a model found compared to the dummy discovery rate achieved by randomly selecting test set crystals, formally the DAF is the ratio of the precision to the prevalence.
-The maximum possible DAF is the inverse of the prevalence or dummy discovery rate which on our dataset is .
-Thus the current SOTA of 3.85 achieved by MACE leaves room for improvement.
-However, evaluating each model on the much smaller subset of its 10k most stable predictions as shown in table 3, CHGNet takes the lead in all 9 metrics, in particular achieving an impressive DAF of 5.29 which approaches optimal performance.
-We find a large performance gap between models that make one-shot predictions directly from unrelaxed inputs (MEGNet, Wrenformer, CGCNN, CGCNN+P, ALIGNN, Voronoi RF) compared to UIPs that learn from force and stress labels to emulate DFT relaxation (CHGNet, M3GNet, MACE).
-While the F1 scores and DAFs of force-free models are surprisingly less affected, their regression metrics such as and RMSE are significantly worse.
-Of the force-free models, only ALIGNN, BOWSR and CGCNN+P achieve positive predictive power .
+Table 1 shows performance metrics for all models included in the initial release of Matbench Discovery reported on the unique protostructure subset.
+EquiformerV2 + DeNS takes the top spot and emerges as the current SOTA for ML-guided materials discovery.
+It outperforms the other models on all 9 reported metrics.
+When computing metrics in the presence of missing values or obviously pathological predictions (error of 5 eV/atom or greater) we assign the dummy regression values and a negative classification prediction to these points.
+The discovery acceleration factor (DAF) measures how many more stable structures a model found compared to the dummy discovery rate achieved by randomly selecting test set crystals.
+Formally the DAF is the ratio of the precision to the prevalence.
+The maximum possible DAF is the inverse of the prevalence which on our dataset is .
+Thus the current SOTA of 5.04 achieved by EquiformerV2 + DeNS leaves room for improvement.
+However, evaluating each model on the subset of 10k materials each model ranks as being most stable (see table S2), we see an impressive DAF of 6.33 for EquiformerV2 + DeNS which is approaching optimal performance for this task.
+We find a large performance gap between models that make energy-only predictions directly from unrelaxed inputs (MEGNet, Wrenformer, CGCNN, CGCNN+P, ALIGNN, Voronoi RF) compared to UIPs that learn from force and stress labels and are used to emulate DFT relaxation before predicting a final energy.
+While the F1 scores and DAFs of energy-only models are surprisingly performant, their regression metrics such as and RMSE are significantly worse.
+Of the energy-only models, only ALIGNN, BOWSR and CGCNN+P achieve positive coefficients of determination, .
 Negative means model predictions explain the observed variation in the data less than simply predicting the test set mean.
@@
 This reveals an important shortcoming of as a metric for classification tasks like stability prediction.
-The results for M3GNet depart from the trend that F1 is broadly rank-correlated with the other classification metrics.
-M3GNet achieves a high true positive rate () but a notably lower true negative rate () than CHGNet and ALIGNN with the closest F1 scores.
-Figure 2 provides a visual explanation for this observation.
-M3GNet has a lower rolling mean of the absolute errors (rolling MAE) as a function of hull distance for materials above the convex hull (see right half of plot) but incurs comparably large errors for materials below the hull (left half of plot).
-Since , lower error for energies above the hull increases both TN and decreases FP, resulting in the high TPR values observed.
 The reason CGCNN+P achieves better regression metrics than CGCNN but is still worse as a classifier becomes apparent from fig.
-7 by noting that the CGCNN+P histogram is more sharply peaked at the 0 hull distance stability threshold.
+S5 by noting that the CGCNN+P histogram is more sharply peaked at the 0 hull distance stability threshold.
 This causes even small errors in the predicted convex hull distance to be large enough to invert a classification.
@@
 This simulates exactly how these models would be used in a prospective materials discovery campaign and reveals how a model’s performance changes as a function of the discovery campaign length.
-As a practitioner, you have a certain amount of resources available to validate model predictions.
-These curves allow you to read off the best model given these conditions and based on the optimal trade-off between fewer false positives (precision) or fewer negatives (recall) for the discovery task at hand.
-In this case, it so happens that CHGNet achieves the highest precision and recall at any number of screened materials.
-A line terminates when a model believes there are no more materials in the WBM test set below the MP convex hull.
+As a practitioner, you typically have a certain amount of resources available to validate model predictions.
+These curves allow you to read off the best model given these constraints.
+For instance, plotting the results in this manner shows that CHGNet initially achieves higher precision, than models such as EquiformerV2+DeNS, ORB MPTrj, SevenNet, and MACE that report higher precision across the whole test set.
+In Figure 2 each line terminates when the model believes there are no more materials in the WBM test set below the MP convex hull.
 The dashed vertical line shows the actual number of stable structures in our test set.
-All models are biased towards stability to some degree as they all overestimate this number, most of all BOWSR by 133%, least of all MEGNet by 30%.
+All models are biased towards stability to some degree as they all overestimate this number, most of all BOWSR by 133%.
 This is only a problem in practice for exhaustive discovery campaigns that validate all stable predictions from a model.
 More frequently, model predictions will be ranked most-to-least stable and validation stops after some pre-determined compute budget is spent, say, 10k DFT relaxations.
-In that case, the concentration of false positive predictions that naturally accumulates near the less stable end of the candidate list can be avoided with no harm to the campaign’s overall discovery rate (see table 3 a similar metrics table considering only the 10k materials predicted by each model to be furthest below the known convex hull).
-The diagonal Optimal Recall line would be achieved if a model never made a false negative prediction and stopped predicting stable crystals exactly when the true number of stable materials is reached.
-Zooming in on the top-left corner of the precision plot, we observe that CHGNet is the only model without a sudden drop in precision right at the start of the discovery campaign.
-This means CHGNet is the only model whose first few hundred most stable materials are largely actually stable.
-MACE in particular suffers from several initial failure cases.
-MACE underpredicts the energy of these unstable materials by several eV/atom (see fig.
-4), resulting in MACE’s initial precision starting at 0 before abruptly climbing to 0.8.
-We experienced this behavior with multiple independently trained MACE variants; even more so with a checkpoint we received directly from the MACE authors which was trained on the M3GNet training set.
-The results shown here are from a superior MACE we trained ourselves on the much larger MPtrj dataset.
-Even M3GNet, the other universal potential and 2nd best model, exhibits the early-on precision drop.
-As a result, CHGNet has a strong lead over M3GNet until reaching 3k screened materials.
-From there, CHGNet and M3GNet slowly converge until they almost tie at a precision of 0.52 after 56k screened materials.
-At that point, CHGNet’s list of stable predictions is exhausted while M3GNet continues, dropping in precision to 0.45 at 76 k, attributable to many false positives near the end of the list of stable predictions.
-All force-free models exhibit a much worse case of early-on precision drop, falling to 0.6 or less in the first 5k screened materials.
+In that case, the concentration of false positive predictions that naturally accumulates near the less stable end of the candidate list can be avoided with no harm to the campaign’s overall discovery rate (see table S2 where even the DAF of the worst performing model benchmarked, Voronoi RF, jumps from 1.58 to 2.49).
+The diagonal ‘Optimal Recall’ line on the recall plot in Figure 2 would be achieved if a model never made a false negative prediction and stopped predicting stable crystals exactly when the true number of stable materials is reached.
+Examining the UIP models, we find that they all achieve similar recall values, ranging from approximately 0.75 to 0.86..
+This is substantially smaller than the variation we see in the precision for the same models 0.44-0.77.
+Inspecting the overlap, we find that the intersection of the models’ correct agreements accounts for a precision of only 0.57 within the 0.75-0.86 range, with just 0.04 of the examples where all models are wrong simultaneously.
+These results indicate that the models are making meaningfully different predictions.
+Examining the precision plot in Figure 2, we observe that the energy-only models exhibit a much more pronouced drop in their precision early-on, falling to 0.6 or less in the first 5k screened materials.
 Many of these models (all except BOWSR, Wrenformer and Voronoi RF) display an interesting hook shape in their cumulative precision, recovering again slightly in the middle of the simulated campaign between 5k and up to 30k before dropping again until the end.
-Figure 2 provides a visual representation of the reliability of different models based on the rolling mean absolute error (MAE) of model-predicted hull distances as a function of DFT distance to the Materials Project (MP) convex hull.
+Figure 3 provides a visual representation of the reliability of different models as a function of a material’s DFT distance to the MP convex hull.
+The lines show the rolling mean absolute error (RMAE) of model-predicted hull distances vs DFT.
 The red-shaded area, which we coin the ‘triangle of peril’, emphasizes the zone where the average model error surpasses the distance to the stability threshold at 0 eV.
-As long as the rolling MAE remains within this triangle, the model is most susceptible to misclassifying structures.
-Because the average error is larger than the distance to the classification threshold at 0, it is large enough to flip a correct classification into an incorrect one (if the error happens to point toward the stability threshold).
-The sooner a model leaves the triangle on the left side, the less likely it is to incorrectly predict stable structures as unstable, thereby reducing false negatives.
-On the right side, early exits result in a lower likelihood of predicting unstable structures as stable, leading to fewer false positives.
-On the left side, CHGNet exits the triangle first followed closely by ALIGNN.
-We can expect fewer false negatives from CHGNet than from other models.
-On the right side, CHGNet and M3GNet exit simultaneously, at a remarkably low error, followed shortly after by MACE at .
-All 3 UIPs therefore produce fewer false positives than other models.
-For highly unstable structures, first CHGNet and later M3GNet are both overtaken by Wrenformer for highly unstable structures towards the right end of the plot.
-Overall, this visualization underscores the importance of considering a model’s average error relative to the distance to the stability threshold when evaluating its performance as a material stability predictor.
-Low overall hull distance error can be a misleading metric if that error is not particularly low at the decision boundary.
-The imbalance between the left and right half of the plot shows that models are more prone to false negative predictions, even for very stable materials far below the known hull, than predicting very unstable materials as stable.
-As the convex hull becomes more thoroughly sampled by future discovery, the fraction of unknown stable structures decreases, naturally leading to less enriched future test sets.
-This has several implications, including allowing for higher maximum DAFs but also skewing the hull distance distribution towards positive values (i.e.
-the right half of fig.
-2) as there are fewer undersampled chemical spaces.
-Consequently, to have accurate stability classifications, model accuracy will need to improve concomitantly.
-For fig.
-2, this means models need to be much more accurate to exit the shaded triangle in the left half of the plot.
+As long as the rolling MAE remains within this triangle, the model is highly susceptible to misclassifying structures.
+The average error in this region is larger than the distance to the classification threshold at 0, consequently in cases where the error points towards the stability threshold it would be large enough to flip a correct classification into an incorrect one.
+The sooner a model exits the triangle on the left side (negative DFT hull distance), the less likely it is to incorrectly predict stable structures as unstable, thereby reducing false negatives.
+Exciting early on the right side (positive DFT hull distance) results in a lower likelihood of predicting unstable structures as stable, leading to fewer false positives.
+All models tend to have lower rolling error towards the plot’s left edge compared to the right edge.
+This imbalance shows that models are more prone to false positive than false negative predictions.
+In other words, all models are less likely to predict a material at DFT hull distance as unstable than a material at DFT hull distance as stable.
+From a practitioner’s standpoint, this is undesirable as attempting to validate an unstable material usually has a much higher opportunity cost than missing a stable one.
+We hypothesize this error imbalance is due to the MP training set having an uncharacteristically high share of stable materials, causing statical models trained on it to be biased towards making low-energy predictions even for high-energy atomic configurations.
+Training on datasets with more high-energy structures, such as Alexandria [schmidt2024improving] and OMat24 [barroso2024open], would be expected to improve performance by balancing out this source of bias.
 4 Discussion
-We have demonstrated the effectiveness of ML-based triage in HT materials discovery and posit that the benefits of including ML in discovery workflows now clearly outweigh the costs.
-LABEL:{tab:metrics-table-uniq-protos} shows in a realistic benchmark scenario that several models achieve a discovery acceleration greater than 2.5 across the whole dataset and up to 5 when considering only the 10k most stable predictions from each model (table 3).
-When starting this project, we were unsure which is the most promising ML methodology for HT discovery.
-Our findings demonstrate a clear superiority in the accuracy and extrapolation performance of UIPs like CHGNet, M3GNet and MACE.
+We have demonstrated the effectiveness of ML-based triage in high-throughput materials discovery and posit that the benefits of including ML in discovery workflows now clearly outweigh the costs.
+LABEL:{tab:metrics-table-uniq-protos} shows that in a realistic benchmark scenario that several models achieve a discovery acceleration greater than 2.5 across the whole dataset and up to 6 when considering only the 10k most stable predictions from each model (table S2).
+When starting this project, we were unsure which is the most promising ML methodology for high-throughput discovery.
+Our findings demonstrate a clear superiority in the accuracy and extrapolation performance of UIPs.
 Modeling forces enables these models to chart a path through atomic configuration space closer to the DFT-relaxed structure from where a more informed final energy prediction is possible.
-The consistently linear log-log learning curves observed in related literature [46] suggest that further decreases in the error of UIPs can be readily unlocked with increased training data.
-Despite its fast pace of progress, the field of universal ML potentials is still in its early stages.
-To realize the full potential of these models, we will need to pivot significant resources to generate large quantities of higher-than-PBE fidelity training data.
-That said, we believe our work (e.g.
-fig.
-8) and others [27] have caught early glimpses that sufficiently large training sets allow us to narrow the performance gap between in-domain and out-of-domain performance.
-This should further increase the discovery acceleration factor, resulting in a virtuous cycle between ML models accelerating the creation of new training data while becoming more accurate and even stronger accelerators after retraining.
-Meanwhile, the setup cost and learning curve of ML potentials are steadily decreasing, likely with a much lower floor than DFT given the intricacies of HT ab-initio simulation.
-Combined with their already strong performance, this suggests that UIPs combined with data accumulation over the next few years may unlock our way to a comprehensive yet cheap map of the PES that truly commoditizes HT materials discovery.
-Despite this optimistic outlook, not only has the predictive power UIPs attained for crystal stability over the last 3 years since [13] sobering 2020 analysis yet to enter common knowledge, but many remaining shortcomings invite further research.
-For example, all models are biased towards stability, overestimating the correct number of stable materials in our test set anywhere between 30% and 133%.
-Moreover, even the best models in our analysis still exhibit a significant fraction of false positive predictions, even for materials as high as above the hull.
-This indicates the need to train future models on even more diverse datasets that are less biased towards ground state crystals than the MP training set we used.
-No such dataset built with MP-compatible DFT settings exists to our knowledge, inviting future efforts in this direction.
-Looking beyond mere thermodynamic stability prediction at zero Kelvin for the purpose of materials discovery, future materials science will also require understanding and predicting the properties under varying environmental conditions such as finite temperature and pressure.
-This is where interatomic potentials can unlock further utility.
-Their force predictions enable insights into dynamical properties.
-One open question of particular relevance is the extent to which such models may aid research into the computational prediction of synthesis pathways.
-Many existing approaches for reaction pathway prediction involve the use of heuristic rules for dealing with the significant added complexity of metastability alongside traditional ground state ab-initio data [47, 48, 49].
-These algorithms stand to benefit massively from more efficient estimates of the reaction energy barriers that future UIPs may provide.
-Provided they reach sufficient accuracy, this may enable timely calculation of approximate reaction barriers [50] and open up a whole new field to high-throughput inquiry.
+As the convex hull becomes more comprehensively sampled through future discoveries, the fraction of unknown stable structures will naturally decline.
+This will lead to less enriched test sets and, consequently, more challenging and discriminative discovery benchmarks.
+However, the discovery task framed here addresses only a limited subset of potential UIP applications.
+We believe that additional benchmarks are essential to effectively guide UIP development.
+These efforts should prioritize task-based evaluation frameworks that address the four critical challenges we identify for narrowing the deployment gap: adopting prospective rather than retrospective benchmarking, tackling relevant targets, using informative metrics, and scalability.
+Looking ahead, the consistently linear log-log learning curves observed in related literature [vonlilienfeld_retrospective_2020] suggest that further decreases in the error of UIPs can be readily unlocked with increased training data.
+This has be borne out in the scaling results of GNoME [merchant2023scaling], MatterSim [yang2024mattersim], Alexandria [schmidt2024improving], and OMat24 [barroso2024open] which all show improvements in performance when training on much larger datasets.
+To realize the full potential of scaling these models, future efforts should deploy their resources to generate large quantities of higher-than-PBE fidelity training data.
+The quality of a UIP model is circumscribed by the quality and level of theory of its training data.
+Beyond simply predicting thermodynamic stability at zero Kelvin, future models will need to understand and predict material properties under varying environmental conditions, such as finite temperature and pressure, to aid in materials discovery.
+In this context, temperature-dependent dynamical properties constitute an area ripe for interatomic potentials.
+Another key open question is how effectively these models can contribute to the computational prediction of synthesis pathways.
+Many existing approaches for reaction pathway prediction involve the use of heuristic rules for dealing with the significant added complexity of metastability alongside traditional ground state ab-initio data [mcdermott_graph-based_2021, aykol_rational_2021, wen_chemical_2023].
+These algorithms will massively benefit from more efficient estimates of reaction energy barriers [yuan2024analytical] and non-crystalline, out-of-equilibrium materials [aykol_thermodynamic_2018], opening up a whole new field to machine learning accelerated inquiry.
 5 Acknowledgments
@@
 Parackal for helping develop the evaluation procedures for Wren upon which this work builds.
-We also thank Rokas Elijosius for assisting in the initial implementation of Wrenformer.
+We also thank Rokas Elijosius for assisting in the initial implementation of Wrenformer and Mark Neumann, Luis Barroso-Luque and Yutack Park for submitting compliant models to the leaderboard.
 We would like to thank Jason Blake Gibson, Shyue Ping Ong, Chi Chen, Tian Xie, Peichen Zhong and Ekin Dogus Cubuk for helpful discussions.
 6 Author Contributions
-Janosh Riebesell: Methodology, Software, Data Curation, Training and Testing Models, Formal Analysis.
-Rhys Goodall: Conceptualization, Software, Formal Analysis.
-Philipp Benner: Software, Training ALIGNN.
-Yuan Chiang: Training MACE.
-Bowen Deng: Curating MPtraj dataset and training CHGNet.
+Janosh Riebesell: Methodology, Software, Data Curation, Investigation (Training: CGCNN, CGCCN+P, Wrenformer, Voronoi RF), Validation, Formal Analysis, Writing – original draft.
+Rhys Goodall: Conceptualization, Software, Validation, Formal Analysis, Writing – original draft, review & editing.
+Philipp Benner: Software, Investigation (Training: ALIGNN, MACE), Writing – original draft.
+Yuan Chiang: Investigation (Training: MACE), Formal Analysis, Writing – review & editing.
+Bowen Deng: Data Curation (MPtrj), Investigation (Training: CHGNet), Writing – review & editing.
+Gerbrand Ceder: Supervision, Funding Acquisition.
+Mark Asta: Supervision, Funding Acquisition.
 Alpha Lee: Supervision.
 Anubhav Jain: Supervision.
-Kristin Persson: Funding Acquisition.
+Kristin Persson: Supervision, Writing – review & editing, Funding Acquisition.
 7 Code availability
@@
 8 Data availability
-We chose the latest Materials Project (MP) [3] database release (v2022.10.28 at time of writing) as the training set and the WBM dataset [29] available at https://figshare.com/articles/dataset/22715158 as the test set for this benchmark.
+We chose the latest Materials Project (MP) [jain_commentary_2013] database release (v2022.10.28 at time of writing) as the training set and the WBM dataset [wang_predicting_2021] available at https://figshare.com/articles/dataset/22715158 as the test set for this benchmark.
 A snapshot of every ionic step including energies, forces, stresses and magnetic moments in the MP database is available at https://figshare.com/articles/dataset/23713842.
 Appendix A Supplementary Information
-A.1 Metrics for 10k materials predicted most stable
-Unlike table 1 which evaluates model performance on the subset of unique WBM structure prototypes, table 2 includes the full WBM test set of 257k materials.
-The 44k additional materials in table 2 comprise discarded earlier due to having a matching structure prototype in MP plus materials which are prototype duplicates of another WBM material with lower energy.
-The most noteworthy difference between the two tables is a drop in DAF for all models in table 2 compared to table 1.
-MACE for example achieves a DAF of 3.5 on the full test set (table 2) compared to 3.85 on the subset of 215.5k materials with unique prototypes (table 1).
-This 10% increase is largely due to a 10% decrease in the fraction of materials below the MP convex hull: 15.3% (32,942 out of 215,488) in table 1 vs 16.7% (42,825 out of 256,963) in the full dataset (table 2).
-Since DAF is the ratio of model’s precision when predicting stability to the prevalence of stable structures, a lower prevalence results in a higher DAF.
-While prototypes are non-trivial to match, thus potentially introducing bias into the test set by trying to deduplicate them, we still opted to feature table 1 in the main text since the removal of overlapping prototypes with MP makes it more closely reflect a model’s true ability to extrapolate to out-of-domain materials.
-Prototypes were matched based on an Aflow-style [51] Wyckoff representation encoding the crystal’s prototype as implemented in aviary.wren.utils.get_aflow_label_from_spglib| [10].
-This string is invariant to relaxation (unless the system changes spacegroup during relaxation, usually by gaining symmetry) and will count structures with different lattice parameters as duplicates if both are expected to relax to the same ground state.
+Note S1.1 Metrics on full test set and for 10k materials predicted most stable
+Unlike table 1 which evaluates model performance on the subset of unique WBM protostructures, table S1 includes the full WBM test set of 257k materials.
+The 44k additional materials in table S1 excluded in table 1 comprise discarded due to having a matching protostructure in MP plus another materials which are protostructure duplicates of another WBM material with lower energy.
+The most noteworthy difference between the two tables is a drop in DAF for all models in table S1 compared to table 1.
+MACE for example achieves a DAF of 3.5 on the full test set (table S1) compared to 3.85 on the subset of 215.5k materials with unique protostructures (table 1).
+This 10% increase is largely due to a 10% decrease in the fraction of materials below the MP convex hull: 15.3% (32,942 out of 215,488) in table 1 vs 16.7% (42,825 out of 256,963) in the full dataset (table S1).
+Since DAF is the ratio of the model’s precision for stability prediction to the prevalence of stable structures, a lower prevalence results in a higher DAF.
+While protostructures are non-trivial to match, thus potentially introducing bias into the test set by trying to deduplicate them, we still opted to feature table 1 in the main text since the removal of overlapping protostructures with MP makes it more closely reflect a model’s true ability to extrapolate to out-of-domain materials.
+Protostructure assignment is based on a combination of Aflow-style prototype labels and the chemical system [hicks_aflow_2021] we use the get_protostructure_label_from_spglib implemented in aviary [goodall2022aviary].
+This string encodes the crystal’s prototype and chemical system and is invariant to symmetry preserving changes to the structure.
+Consequently structures with different lattice parameters are flagged as protostructure duplicates as they would both be expected to relax to the same ground state in experimental conditions.
 A real-world discovery campaign is unlikely to validate all stable predictions from a given model as we did in LABEL:{tab:metrics-table-uniq-protos}.
@@
 Assuming that increases in compute resources will allow average future discovery campaigns to grow in scope, we believe 10k model validations to be a reasonable scope for average campaigns.
-This is what table 3 simulates by calculating classification and regression metrics for the 10k test set materials predicted to be most stable by each model.
+This is what table S2 simulates by calculating classification and regression metrics for the 10k test set materials predicted to be most stable by each model.
 We again show dummy performance in the bottom row.
@@
 However, the bottom row still shows dummy performance across the whole dataset.
-CHGNet and M3GNet achieve an impressive precision of 0.83 and 0.8, respectively.
-In concrete terms, this means in a discovery campaign that validates 10 k model predictions from a search pool of 257 k crystals that are chemically dissimilar from the training set and of which 0.167 are stable, CHGNet and M3GNet would deliver 4 stable structures for every 5 predictions validated.
-We emphasize that in light of the significant resulting increase in stability hit rate, these models are well worth integrating into future materials searches.
-A.2 ROC Curves
+Note S1.2 ROC Curves
 A material is classified as stable if the predicted is below the stability threshold.
-Since all models predict (and M3GNet predicted energies are converted to formation energy before stability classification), they are insensitive to changes in the threshold .
+Since all models predict , they are insensitive to changes in the threshold .
 The receiver operating characteristic (ROC) curve for each model is plotted in fig.
-3.
-M3GNet wins in the area under curve (AUC) with 0.87, coming in 1.34x higher than the worst model, the Voronoi Random Forest.
+S1.
 The diagonal ‘No skill’ line shows the performance of a dummy model that randomly ranks material stability.
-A.3 Parity Plots
-Figure 4 shows that all models do well for materials far below the convex hull (left side of the plot).
+Note S1.3 Parity Plots
+Figure S2 shows that all models do well for materials far below the convex hull (left side of the plot).
 Performance for materials far above the convex hull is more varied with occasional underpredictions of the energy of materials far above the convex hull (right side).
@@
 Wrenformer is the only model with a large number of severe energy overpredictions at along the positive axis.
-We investigate these failure cases in more detail in fig.
-5 and find these overpredictions to be dominated by spacegroup 71 with poor representation in the training data.
+We investigated these failure cases in more detail and found these overpredictions to be dominated by spacegroup 71 with poor representation in the training data.
+Digging into the MP-crystals-2022.10.28 dataset we see that on post analysis there are a large number of A2BC compounds in spacegroup 71 in the v2022.10.28 that are trapped in local minima and thus much higher in energy.
+Given that Wrenformer nominally predicts the energy the global minima conditioned on the protostructure these are unsuitable for training Wrenformer and their inclusion leads to the errors seen on similar, correctly relaxed structures, in the WBM test set.
 The other anomalous model is MACE with several severe underpredictions at along the negative axis.
@@
 Beyond these MACE outliers visible in the plot, MACE exhibits another rare but reproducible type of failure case, in which the final predicted energy after relaxation is off by several orders of magnitude.
-The largest ‘derailed’ prediction was eV/atom for wbm-3-31970 (formula HIr).
+The largest ‘derailed’ prediction was eV/atom for wbm-3-31970 (formula H2Ir).
 In each case, the MACE relaxation exhausted the maximum number of ionic steps set to 500 and caused a volume implosion from initial cell volumes of hundreds to relaxed cell volumes of tens of cubic Angstrom.
@@
 Since these derailed values are easily identified in practice when actually performing a prospective discovery campaign, we excluded them from the MACE parity plat and all other downstream analyses.
-A.4 Spacegroup prevalence in Wrenformer failure cases
-Figure 5 shows 456 + 194 (70%) of the failure cases in the shaded rectangle are two prototypes in spacegroup 71.
-The occurrence of those same prototypes in the MP training set shows almost no data support for the failing prototypes.
-This suggests the reason Wrenformer fails so spectacularly on these structures is that it cannot deal with structure prototypes it has not seen at least several hundred examples of in its training data.
-This suggests that there are stronger limitations on how much the discrete Wyckoff-based representation can extrapolate to new prototypes compared to the smooth local-environment-based inputs to GNN-type models.
-A.5 Hull Distance Box plot
+Note S1.4 Hull Distance Box plot
+Figure S4 shows a box-plot of the errors in predicting the distance to the convex hull.
+The box demarks the interquartile range and the whiskers show the 5th and 95th percentiles.
 BOWSR has the largest median error, while Voronoi RF has the largest IQR.
-Note that MEGNet and CGCNN are the only models with a positive median.
-Their hull distance errors are biased towards more frequently predicting thermodynamic instability, explaining why they are closest to correctly predicting the overall number of stable structures in the test set (see fig.
-1).
-A.6 Classification Histograms using Model-Predicted Energies
-Note the CGCNN+P histogram is more strongly peaked than CGCNN’s which agrees better with the actual DFT ground truth distribution of hull distances in our test set.
+Looking at the models we see that there is a systematic trend to underpredict the distance to the convex hull.
+This trend can be rationalized give the bias in the Materials Project towards stable and therefore low formation energy structures.
+In contrast, due to it’s construction the WBM test set is distributionally distinct with a higher average formation energy.
+Therefore we propose that models trained on the Materials Project are likely result in such systematic underpredictions and this affect should be able to be addressed by training on additional data from a similar distribution to WBM i.e.
+Alexandria [schmidt2024improving].
+Note S1.5 Classification Histograms using Model-Predicted Energies
+Figure S5 shows histograms of the model predicted distances to the convex hull colored by the stability classification.
+These plots allow for practitioners to assess how the accuracy varies for materials depending on how far above or below the hull they are predicted to lie.
+The most pronounced affect we see from the histograms is that the CGCNN+P histogram is more strongly peaked than CGCNN’s giving much better agreement with the actual DFT ground truth distribution of hull distances for the test set.
 This explains why CGCNN+P performs better as a regressor, but also reveals how it can perform simultaneously worse as a classifier.
 By moving predictions closer to the stability threshold at above the hull, even small errors are significant enough to tip a classification over the threshold.
-A.7 Measuring extrapolation performance from WBM batch robustness
+Note S1.6 Measuring extrapolation performance from WBM batch robustness
 As a reminder, the WBM test set was generated in 5 successive batches, each step applying another element replacement to an MP source structure or a new stable crystal generated in one of the previous replacement rounds.
@@
 We observe this effect for some models much more than others.
-Figure 8 shows the rolling MAE as a function of distance to the convex hull for each of the 5 WBM rounds of elemental substitution.
-These plots show a stronger performance decrease for Wrenformer and Voronoi RF than for UIPs like M3GNet, CHGNet, MACE and even force-less GNNs with larger errors like MEGNet and CGCNN.
+Figure S6 shows the rolling MAE as a function of distance to the convex hull for each of the 5 WBM rounds of elemental substitution.
+These plots show a stronger performance decrease for Wrenformer and Voronoi RF than for UIPs and even force-less GNNs with larger errors like ALIGNN, MEGNet and CGCNN.
+Figure S6 shows the rolling MAE for different models split out by calculation batch in the WBM data set.
 MEGNet and CGCNN both incur a higher rolling MAE than Wrenformer across all 5 batches and across most or all of the hull distance range visible in these plots.
@@
 However, the loss of predictive forces and stresses may make Wrenformer unattractive for certain applications even then.
-A.8 Largest Errors vs. DFT Hull Distance
+Note S1.7 Largest Errors vs. DFT Hull Distance
 Given the large variety of models tested, we asked whether any additional insight into the errors can be gained from looking at how the predictions vary between different models.
 In fig.
-9 we see two distinct groupings emerge when looking at the 200 structures with the largest errors.
+S7 we see two distinct groupings emerge when looking at the 200 structures with the largest errors.
 This clustering is particularly apparent when points are colored by model disagreement.
-A.9 MEGNet formation energies from UIP-relaxed structures
-Interestingly, table 4 shows that MEGNet RS2RE performs worse than CHGNet and M3GNet even when it has access to the ground truth DFT relaxed structures when making relaxed energy predictions.
-While the regression and classification metrics for MEGNet RS2RE are better than M3GNet→MEGNet and CHGNet→MEGNet, the difference is surprisingly small, especially compared to the regression performance of standalone MEGNet.
-This suggests that for downstream use, at least by an imprecise method such as another ML model, the UIP-relaxed structures are almost as helpful as the DFT-relaxed ones.
-CHGNet, M3GNet and MEGNet were all trained on slightly different targets.
-While MEGNet was trained to predict formation energies, M3GNet and CHGNet were trained on raw DFT energies with the important distinction that CHGNet targets include MP2020 energy correction scheme [52] while M3GNet targets do not.
-Changing the target from raw DFT energies to formation energies in principle merely constitutes a change of gauge since the difference is just a linear transformation of subtracting the elemental reference energies weighted by composition.
-But in practice, one might expect the problem of stability prediction to be easier when trained on formation energies, as the model output is one less step removed from the final target of convex hull distance.
-However, empirical evidence so far suggests the opposite.
-Both UIP→MEGNet combos perform worse than those same UIPs by themselves.
-There are too many confounding effects at play to draw firm conclusions but this tentatively suggests using formation energy as targets offers no benefits over raw DFT.
-A.10 Exploratory Data Analysis
+Note S1.8 Exploratory Data Analysis
 To give high-level insights into the MP training and WBM test set used in this work, we include element distributions for structures in both datasets (fig.
-11).
+S8).
 To show how frame selection from MP structure relaxation affected relative elemental abundance between MP relaxed structures and the snapshots in MPtrj, fig.
-12 shows element occurrence ratios between MPtrj and MP.
-Believing MPtrj to be an influential dataset for the near-term continued development of universal interatomic potentials, we plot histograms showing the distributions of target values for energies, forces, stresses and magnetic moments in fig.
-14.
+S9 shows element occurrence ratios between MPtrj and MP.
 Similarly, fig.
-15 plots the distribution of formation energies and convex hull distances (with respect to the convex hull spanned by MP materials only) of the WBM test set.
+S10 shows the elements-per-structure distribution of MP, MPtrj and WBM, normalized by dataset size.
+The mode of all three datasets is 3, but WBM’s share of ternary phases is noticeably more peaked than MP’s, which includes small numbers of unary and senary phases.
+fig.
+S11 plots the distributions of formation energies, decomposition energies and convex hull distances (with respect to the convex hull spanned by MP materials only) for the MP and WBM data sets.
 This plot not only gives insight into the nature of the dataset but also emphasizes the increased difficulty of stability vs formation energy prediction arising from the much narrower distribution of convex hull distances compared to the more spread-out formation energy distribution.
-To get a visual sense of coverage of atomic configuration space afforded by WBM compared to MP (as an indicator for the degree of out-of-distribution extrapolation required), we project matminer [53] features for all initial structures as well as relaxed MP structures onto two UMAP dimensions in fig.
-10.
-Finally, fig.
-12(a) shows the elements-per-structure distribution of MP, MPtrj and WBM, normalized by dataset size.
-The mode of all three datasets is 3, but WBM’s share of ternary phases is noticeably more peaked than MP’s, which includes small numbers of unary and senary phases.
+For WBM we see that the formation energy distribution exhibits much wider spread than the convex hull distance distribution, spanning almost vs less than spread in the hull distances.
+This highlights why stability prediction is a much more challenging task than predicting energy of formation.
+It requires correctly ranking the subtle energy differences between chemically similar compounds in the same chemical system rather than comparing a single material with the reference energies of its constituent elements.
+DFT has been shown to significantly benefit from the systematic cancellation of errors between chemically similar systems when trying to identify the lowest-lying polymorph [hautier_accuracy_2012].
+This beneficial cancellation has yet to be conclusively demonstrated for ML stability predictions.
+So far, only the lack thereof has been shown in [bartel_critical_2020] where they encountered a much more random error distribution among similar chemistries than simulations from first principles.
+Finally, Believing MPtrj to be an influential dataset for the near-term continued development of universal interatomic potentials, we plot histograms showing the distributions of target values for energies, forces, stresses and magnetic moments in fig.
+S12.
 The histogram in fig.
-12(b) shows the distribution of the number of sites in MPtrj structures.
+12(e) shows the distribution of the number of sites in MPtrj structures.
 The inset displays the same histogram log-scaled y-axis as well as a cumulative line to show that 90% of MPtrj structures contain fewer than 70 sites.
-A.11 Predicted Formation Energy Analysis
-We show the distribution for all materials in the WBM test set of model-predicted formation energies vs DFT ground truth in fig.
-17 and the distribution of convex hull distance errors projected onto elements by composition for MACE and CHGNet in fig.
-18.
+fig.
+S13 shows the forces and magnetic moments for the MPTrj dataset projected onto different elements.